### v2

# Unsupervised and Source-Free Ranking of Biomedical Segmentation Models

Abstract
Model transfer presents a solution to the challenges of segmentation in the biomedical community, where the immense cost of data annotation is a major bottleneck in the use of deep learning. At the same time, hundreds of models get trained on biomedical data, submitted to challenges, and posted in model zoos and repositories. A major hurdle to wider adoption of pre-trained models lies in the lack of methods for best model selection. While such methods have been proposed for classification models, semantic and instance segmentation model ranking remain largely unaddressed, especially in a practically important setting where no labels are available on the target dataset. Similarly, if unsupervised domain adaptation is used, practitioners are faced with the task of selecting the best adapted model without target domain labels. Building on previous work linking model generalisation and consistency under perturbation, we propose the first unsupervised and source-free transferability estimator for semantic and instance segmentation tasks. We evaluate on multiple segmentation problems across biomedical imaging, finding a strong correlation between the rankings based on our estimator and rankings based on target dataset performance.
1 Introduction
Segmentation is a ubiquitous problem in biomedical image analysis as it plays a key role in the interpretation of biological and anatomical structures. While modern deep learning methods have significantly improved segmentation accuracy, their practical use remains limited by the need for labour-intensive, pixel-wise annotation of training data. Transfer learning [75, 45] aims to address this problem through re-use of networks trained on other datasets (“source”), either similar to the dataset of current interest (“target”) or sufficiently large to sample all possible targets. The latter approach has been prevalent in the analysis of natural images, while in biomedical imaging, many practitioners rely on training domain-specific models on their own data [28]. Community efforts such as the BioImage Model Zoo (BMZ) [49] make many such models publicly available, public challenges like the ToothFairy series [9] drive the creation and sharing of new models, and emerging large generalist models have had their first successes in the biomedical domain [3, 51, 64]. These initiatives raise challenges in selecting the best model for a target dataset when multiple models—trained on the same task but with varying settings, architectures, or source data—are available.
For the task of image classification, multiple supervised model selection/transferability ranking methods have been introduced [57, 16, 27, 80, 47, 74, 7, 36]. It has also been suggested that transferability ranking for semantic segmentation can be achieved as a simple extension of classification approaches [1]. In our experiments, this suggestion does not hold for biomedical imaging data, where classification-derived transferability metrics fail to reliably correlate with model performance on the target datasets. Furthermore, while semantic segmentation can be expressed as a pixel classification problem and thus derived from classification approaches, instance segmentation cannot. Instance segmentation outputs an unordered, variable‑length set of instances (masks, boxes) that only become discrete after model‑ and hyperparameter‑dependent instantiation steps (e.g., thresholding, binarisation, NMS/post‑processing). Instantiation results in a permutation invariant set of masks that have no stable, fixed‑dimensional correspondence between feature vectors, predictions, and target labels for the metrics to use. To the best of our knowledge, the problem of instance segmentation transferability ranking thus remains unaddressed, both for natural and for biomedical images. Supervised ranking exploiting target domain labels has other downsides: annotation, especially in 3D, can take weeks of expert time without any guarantee that a representative sample has been annotated. Unsupervised ranking is, therefore, much more practically useful. Also, recent advances in Unsupervised Domain Adaptation (UDA) have enabled fine-tuning without target domain labels, bringing forward the related problem of unsupervised evaluation of adapted model performance. First such methods have been introduced [59, 44, 78], again largely for classification models. However, in our experiments, the state-of-the-art for unsupervised validation [78] consistently underperforms for semantic segmentation. Also, like all methods directly extended from classification tasks, they are not applicable to instance segmentation.
Paper Contribution. The aim of this paper is to address the challenge of unsupervised segmentation model ranking to facilitate wider adoption of deep learning-based segmentation in the biomedical domain, where the difficulty of the annotation process and resource limitations hinder supervised training or model selection. Additionally, to ensure practical applicability within model repositories, we aim to operate in a source-free setting.
Building on the approaches proposed for model generalisation and uncertainty estimation, we (i) introduce a novel zero-shot transferability estimator for semantic and instance segmentation problems (Fig. 1). The estimator is based on prediction consistency under feature or input space perturbation and is both source-free and unsupervised. We (ii) evaluate the estimator on semantic and instance segmentation problems that represent highly popular tasks in microscopy: cell and nucleus segmentation in light microscopy and mitochondria segmentation in electron microscopy. We compare a diverse range of architectures and sizes, from small domain-specific models to generalist segmentation tools. Additionally, we (iii)
apply the estimator to models submitted to the ToothFairy challenge series [9], a benchmark for multi‑class 3D CBCT (Cone Beam Computed Tomography) semantic segmentation, reproducing challenge average DICE ranking without access to ground-truth. (iv) Beyond our main task of zero-shot transferability ranking, we apply our approach to rank models after UDA. (v) An extensive set of experiments shows that model rankings produced by our estimator are strongly correlated with target dataset performance, providing a simple but powerful model selection method enabling effective reuse of pre-trained models for practitioners in the field. Code will be made publicly available on acceptance.
2 Related Work
Qualitative approaches comparing datasets are common practice, but unreliable, as imperceptible differences in source and target dataset distributions, model architectures, and training regimes impact likely transfer success [62, 42, 45]. In the following, we review the related work in supervised transferability prediction and post-UDA model selection, highlighting work from the domain generalisation and uncertainty estimation fields that provided inspiration.
Supervised Transferability Metrics. Transferability metrics fall into two main categories: (1) label-comparison–based methods, such as LEEP [47], and (2) feature-embedding–based methods, which analyze the embedding space on target data with guidance from target labels. Examples include: LEEP [36], LogME [80], H-score [7], Regularised H-score [27], GBC [52], NCTI [74], and CC-FV [79], which, unlike the others, is explicitly designed for semantic segmentation. The above approaches require target domain labels and were mostly designed for classification. Although our proposed method is unsupervised, in the absence of unsupervised baselines, we select all these models as baselines for our experiments.
Unsupervised Validation. Many popular UDA methods rely on self or adversarial training [32, 12] and are notoriously unstable, occasionally even leading to model degeneration [71, 73]. The domain of UDA validation, therefore, addresses a problem close to ours: how to select the best adapted model without target domain labels. The state-of-the-art in this field is the Transfer Score (TS) [78], developed for classification, but also applied to segmentation. It measures the classifier bias of a UDA model in the target domain along with the feature space transferability and discriminability. We use this score as a baseline in our experiments.
Domain Generalisation and Uncertainty Estimation. The task of model generalisation evaluation aims to train models that generalise better. While our goal is to estimate, rather than improve transferability, we can exploit similar approaches to measuring the generalisability, assuming that more generalisable models also transfer better. Here, consistency-based approaches [15, 61, 2] potentially satisfy our requirements for an unsupervised estimator and, although introduced for classification, could be extended to serve as a basis for semantic segmentation model ranking.
Early works [2, 61] suggest that prediction invariance under test-time-augmentation can be linked to model generalisability. However, they only focus on the in-domain (source) setting. In the case of [2] they rely on a carefully weighted penalty score to evaluate prediction invariance and [61] uses ground-truth to directly measure changes in the classification accuracy. Effective Invariance (EI) [15] was the first to address out-of-distribution (OOD) generalisability. EI measures both the consistency and confidence of probabilistic model outputs, showing that a combination of both is necessary for classification. However, it makes the method susceptible to failure for poorly calibrated models, which is especially relevant in the OOD transfer setting [50]. Furthermore, the EI cannot be extended to instance segmentation.
Consistency to perturbations has also been widely investigated in the uncertainty estimation field, where variance of prediction can be used as a measure of epistemic uncertainty [19, 39, 33, 43]. Uncertainty estimation aims to find well calibrated estimates of model confidence that correlate with true performance. The field is not concerned with transferability or generalisability of models and often makes use of supervised target labels to help calibrate uncertainty. Most relevant to our task are the post-hoc uncertainty estimators such as [39, 33, 43, 70] as ranking cannot influence model training. Rich theory links these approaches to Bayesian Neural Networks and Monte-Carlo Dropout [19], providing motivation for probing model responses under both input and feature space perturbations. As our ideal transferability estimator also needs to strongly correlate with model performance on target data, the work in these two domains provides inspiration for our approach.
3 Methods
3.1 Problem Definition
For unsupervised model transfer, a pre-trained model , trained on a supervised source domain is applied to an unlabelled target dataset taken from a target domain . Inference on can either be performed directly or after UDA, which further trains on an unlabelled subset of . We assume a transductive transfer setting [20], where the marginal distributions of the domains differ, , resulting in a domain shift, , but the source and target task are fixed (e.g., nuclei segmentation). We introduce an unsupervised consistency-based transferability metric (CTE) to correlate with the true performance (e.g., F1 score), which is not available without GT. Given a set of candidate source models , the goal of a transferability metric is to preserve the ranking if and only if .
3.2 Motivation
We propose to estimate model transferability based on the consistency of model outputs under perturbation. Focusing the analysis solely on model outputs enables a fully unsupervised and source-free transfer ranking. Since the output spaces of candidate models are fixed and directly comparable—unlike their feature spaces, which may differ substantially in composition and dimensionality across architectures—our approach provides a uniform basis for evaluation. Moreover, by operating in the final segmentation output space, our method allows, for the first time, direct assessment of the transferability of segmentation models, even across varied prediction processing/instantiation approaches.
The motivation for using consistency as a measure of transferability arises from the following considerations: a source model trained on a specific data distribution is expected to learn decision boundaries that partition this space by class. When applied to a target domain with a different distribution, these boundaries may fail to separate target classes effectively or may intersect clusters of target pixel embeddings. To assess transferability, we introduce small perturbations—either to the model input or its features—and observe whether pixel embeddings cross decision boundaries in the output prediction. Measuring such pixel class flips provides an assumption-free estimate of transferability, in contrast to prior methods that rely on the assumptions of feature-space compactness or separability, often modelled with Gaussian distributions [52, 36]. This is advantageous for semantic/instance segmentation, where common losses (e.g., combinations of Dice and cross-entropy) do not explicitly enforce feature compactness.
Prediction consistency can be viewed as a proxy for the margin of a model’s decision boundaries with respect to the target data. Models that transfer well are expected to exhibit larger margins, leading to better-separated classes and predictions that are less sensitive to perturbations. This intuition aligns with prior discussions on manifold smoothness and margin distribution [29, 46]. Related reasoning is also adopted in GBC [52], which uses class feature separability as a transferability metric. This metric, however, requires access to target labels, so we choose to probe for perturbation-induced class changes instead.
From this perspective, using input image perturbations, as customary in domain generalisation, or feature space perturbations, popular in uncertainty estimation, are conceptually equivalent. Unlike input perturbations for data augmentation in training, consistency perturbations need not represent realistic image changes between source and target distributions. Their purpose is to perturb pixel embeddings, testing the suitability of a model’s decision boundary for target data by measuring the number of output class label flips. Also, feature space perturbations do not necessarily correspond to realistic input image changes. In our experiments (Sec. 4), we evaluate several perturbation strategies and find all are effective.
3.3 Consistency Transferability Estimator (CTE)
Considering a single target dataset , we rank the transfer performance of a set of models on by measuring pixelwise prediction consistency between perturbed and unperturbed predictions (Fig. 1). For classification or semantic segmentation, we can directly compare class labels of the pixels.
Firstly, we propose to extend the Effective Invariance (EI) [15], originally formulated for classification tasks, to semantic segmentation. EI measures the invariance of a single prediction to test-time perturbations, which for segmentation can be formulated as:
where and represent the output of a model after softmax and its thresholded value, with representing the perturbed outputs. We refer to the CTE-EI as a “soft” metric as it measures both the consistency and confidence of the probabilistic output. This introduces a reliance on good model calibration which is not guaranteed in transfer [50]. We therefore propose another consistency measure based on the thresholded model output (“hard” consistency). This measure directly compares the number of pixelwise changes between the perturbed and unperturbed predictions, using the Hamming distance normalised by the number of pixels:
This is equivalent to the between the predictions, but considering in terms of pixel flips adds intuition. To counteract the high class imbalance of biomedical data and to allow for multiclass problems, we calculate per class consistency and then apply the weighted average. This is achieved by limiting the pixelwise iterator to the pairwise union of perturbed and unperturbed prediction masks for this class.
For instance segmentation, direct pixel comparison is impossible as instance labels are arbitrary. We therefore modify our approach to measure the consistency of assignment of pixel pairs to the same or different objects, as measured by the Rand Index [56, 67]. To account for class imbalance, we utilise the foreground-restricted Adapted Rand Score (ARS), defined by [4] as:
where is the probability that a pixel has the same label in the perturbed prediction and in the unperturbed prediction, is the probability that a pixel has label in , and is the probability that a pixel has label in .
The consistency score for a single transfer is then calculated as:
where is calculated using one of the above metrics (CTE-EI, NHD, ARS). The median is used for robustness with small dataset sizes. The final ranking within is the descending order of .
3.4 Model Perturbation
Perturbations can be applied either to the model input through test-time augmentations or directly within the feature space, assuming access to intermediate model representations. Both approaches are motivated by prior work in uncertainty estimation [43, 39] and consistency training [48].
We assess prediction robustness to perturbations under transfer to OOD data, thereby probing the suitability of the embedding space and decision boundaries for target embeddings. This differs from evaluating the model’s learned invariance to perturbations within the source domain (e.g., due to data augmentation during training). Even for models trained with augmentations, perturbations can still expose informative variations on target data. To examine this, we rank models trained both with and without augmentation and consider perturbations in both input and feature spaces.
Perturbations must remain tolerable [43]: if they are too strong, even highly transferable models become inconsistent, while overly weak perturbations leave all models perfectly stable. Empirically, we find that a broad range of perturbation magnitudes satisfy this condition allowing effective transfer estimation via CTE (see Supp.).
Input Space Perturbations. As input perturbations, we explored several popular image transformations: additive Gaussian noise, gamma correction, and changes in brightness and contrast. In our experiments we found similar performance across all perturbation types. For brevity we report Gaussian noise, where controls the strength of the perturbation, except for 3D segmentation where we also report gamma results. See Supp. and Tab. 2 for additional experiments.
Feature Perturbation.
Given access to a model, perturbations can be applied directly to intermediate feature layers, while keeping the network weights frozen (greybox model). Motivated by findings in semi-supervised learning [48] and also by work in uncertainty estimation [43], we investigate test-time spatial dropout [66]. Specifically, we apply dropout (1) across all layers of the network, (2) only at the bottleneck, where information is the most redundant [25] (Tabs. 1 and 4) or (3) at the bottleneck and in the skip connection layers (Tab. 2). The dropout rate is controlled by . For a network with layers and dropout perturbation applied at layer the prediction is defined as:
4 Experimental Setting
4.1 Datasets
We evaluate our approach on a range of public segmentation datasets, covering six distinct tasks and multiple modalities. For each transfer prediction experiment, we fix one target dataset and evaluate the transfers from all task-specific models (see Supp. for further details).
Electron Microscopy (EM), Mitochondria. We use four semantic segmentation datasets with neural tissues from different species and EM modalities: [40, 17, 54]. We also reformulated these datasets as a per-patch binary classification task, labelling patches as positive if mitochondria are present. This auxiliary task allows us to disentangle domain shift effects from the intrinsic suitability of classification-based ranking methods for semantic segmentation.
Light Microscopy (LM), Nuclei. We use nine fluorescent nuclei datasets for instance and semantic segmentation taken from a variety of species and LM modalities: [68, 69, 31, 23, 38, 11, 5, 14, 18].
Light Microscopy, Cells. We consider four dense instance cell segmentation datasets, taken from fruit fly, Arabidopsis thaliana and human samples: [18, 77, 76, 53].
Cone-Beam Computed Tomography (CBCT), ToothFairy2. We use the ToothFairy2 datasets taken from a MICCAI2024 challenge [9, 10], comprising 3D CBCT volumes of the human jaw with 42 labelled classes, training the models on the official challenge training set and assessing transferability on the test set acquired from another institution.
4.2 Models
We trained a diverse set of model architectures for transfer comparison. For classification: ResNet18 [24], ResNet50 [24], DenseNet121 [26], MobileNetV2 [60], MobileNetV3 [55]. For binary semantic and instance segmentation: 2D U-Net [58], Residual U-Net [34] and Unetr [22]. We also ranked publicly available models such as [3], Cellpose-SAM [51], and a model from the BMZ [49], id: 5092850 ‘powerful-chipmunk‘ [53]. For multi-class semantic segmentation, we selected top-models from the ToothFairy2 challenge [72] (Isensee et al. and Wang et al.) and further included architectures covering CNNs (2D nnU-Net, residual-encoder 3D nnU-Net [28]), Transformers (SwinUNETR [21]) and State Space Models (UMamba [41], VMamba [37]).
| Task | Unsupervised | Supervised | |||||||||||||||||||
TS | CCFV | LEEP | GBC | Hscore | LEEP | LogME | NCTI | RegHscore | |||||||||
| Mito | K | 0.77 ±0.1 | 0.71 ±0.1 | 0.73 ±0.1 | 0.69 ±0.1 | 0.25 ±0.2 | -0.11 ±0.1 | 0.41 ±0.4 | 0.39 ±0.1 | 0.23 ±0.2 | 0.89 ±0.1 | 0.08 ±0.2 | 0.21 ±0.2 | 0.24 ±0.2 | |||||||
| S | 0.85 ±0.1 | 0.83 ±0.1 | 0.86 ±0.1 | 0.85 ±0.1 | 0.30 ±0.3 | -0.16 ±0.1 | 0.49 ±0.5 | 0.52 ±0.1 | 0.33 ±0.4 | 0.96 ±0.0 | 0.10 ±0.3 | 0.28 ±0.2 | 0.34 ±0.4 | ||||||||
| P | 0.83 ±0.1 | 0.79 ±0.1 | 0.86 ±0.1 | 0.84 ±0.1 | 0.23 ±0.3 | -0.18 ±0.1 | 0.39 ±0.3 | 0.47 ±0.2 | 0.37 ±0.3 | 0.97 ±0.0 | 0.06 ±0.3 | 0.37 ±0.2 | 0.37 ±0.2 | ||||||||
| Nuclei | K | 0.62 ±0.1 | 0.69 ±0.1 | 0.74 ±0.2 | 0.71 ±0.2 | 0.02 ±0.3 | -0.03 ±0.3 | 0.11 ±0.1 | -0.13 ±0.3 | 0.29 ±0.5 | 0.67 ±0.4 | -0.20 ±0.3 | 0.29 ±0.5 | 0.35 ±0.4 | |||||||
| S | 0.77 ±0.1 | 0.82 ±0.1 | 0.85 ±0.1 | 0.84 ±0.1 | 0.10 ±0.4 | -0.05 ±0.4 | 0.12 ±0.1 | -0.16 ±0.4 | 0.25 ±0.6 | 0.77 ±0.3 | -0.28 ±0.3 | 0.25 ±0.6 | 0.36 ±0.4 | ||||||||
| P | 0.98 ±0.0 | 0.97 ±0.0 | 0.90 ±0.1 | 0.62 ±0.4 | -0.09 ±0.4 | -0.16 ±0.5 | 0.32 ±0.2 | -0.08 ±0.4 | 0.31 ±0.4 | 0.94 ±0.0 | -0.49 ±0.3 | 0.30 ±0.3 | 0.34 ±0.3 | ||||||||
4.3 Baselines
We include supervised transfer metrics: LEEP [36], GBC [52], HScore [7], LEEP [47], LogME [80], NCTI [74], Regularised-HScore [27], CC-FV [79], which represent the state-of-the-art for biomedical segmentation. We also use the SOTA for unsupervised validation Transfer Score (TS) [78]. No baselines are available for instance segmentation.
4.4 Performance Evaluation
For all transferability metrics, we calculate the correlation between the metric ranking and the performance score ranking. As is standard [1] we evaluate the correlation scores using: Kendall tau [30], Pearson correlation coefficient [8], and Spearman’s rank correlation coefficient [63]. The metrics vary in the range with indicating perfect positive correlation. The metrics are complementary, indicating if the agreement is influenced more by overall rank differences or local pairwise inconsistencies. While a perfect should be monotonically related to the , linear correlation is not required for transferability metrics.
5 Results
5.1 Semantic Segmentation
Previous work [1] has suggested that supervised transferability metrics developed for classification tasks can be readily extended to semantic segmentation by treating it as pixelwise classification. The standard approach samples a class-balanced set of pixels from each image and applies the transfer metric directly to this per-pixel classification set. However, for our biomedical datasets, this assumption does not hold: existing transfer metrics, even with supervision, fail to predict either direct transfer performance (Tab. 1) or post–fine-tuning performance (see Supp.). To determine if this limitation arises from the biomedical imaging domain [13] or the shift to semantic segmentation, we additionally evaluated these metrics on mitochondria classification.
Fig. 2 shows the performance score (F1) plotted against each transfer score for a single target dataset (Hmito) for both classification and semantic segmentation tasks. The metrics perform well for classification (violet crosses) but, except for LEEP and our CTE, fail to reliably correlate with semantic segmentation rankings (green circles).
| Class | CTE-EI (Gauss) | CTE-NHD (Gauss) | CTE-EI (DropOut) | CTE-NHD (DropOut) | CTE-EI (Gamma) | CTE-NHD (Gamma) | |||||||
| IACs | K | 0.90 | (**) | 0.81 | (**) | 0.87 | (*) | 0.87 | (*) | 1.00 | (**) | 0.90 | (*) |
| S | 0.96 | (**) | 0.93 | (*) | 0.94 | (**) | 0.94 | (**) | 1.00 | (**) | 0.96 | (**) | |
| P | 0.99 | (**) | 0.98 | (**) | 0.83 | (*) | 0.80 | (0.1) | 0.99 | (**) | 0.98 | (**) | |
| Teeth | K | 0.81 | (**) | 1.00 | (**) | 0.67 | (0.1) | 0.75 | (0.1) | 0.71 | (*) | 0.90 | (**) |
| S | 0.89 | (*) | 1.00 | (**) | 0.77 | (0.1) | 0.77 | (0.1) | 0.86 | (*) | 0.96 | (**) | |
| P | 0.95 | (**) | 0.98 | (**) | 0.85 | (**) | 0.94 | (**) | 0.89 | (**) | 0.95 | (**) | |
| Mand. | K | 0.62 | (*) | 0.52 | (0.2) | 0.87 | (*) | 0.87 | (**) | 0.43 | (0.3) | 0.81 | (**) |
| S | 0.82 | (*) | 0.68 | (0.1) | 0.87 | (*) | 0.94 | (*) | 0.54 | (0.2) | 0.89 | (**) | |
| P | 0.73 | (0.1) | 0.65 | (0.1) | 0.89 | (**) | 0.96 | (**) | 0.74 | (0.1) | 0.80 | (*) | |
| Sinus. | K | 0.43 | (0.3) | 0.81 | (*) | 0.87 | (*) | 0.83 | (*) | 0.43 | (0.2) | 0.90 | (**) |
| S | 0.50 | (0.3) | 0.89 | (*) | 0.84 | (*) | 0.94 | (*) | 0.57 | (0.2) | 0.96 | (**) | |
| P | 0.38 | (0.4) | 0.73 | (0.1) | 0.88 | (**) | 0.90 | (**) | 0.38 | (0.4) | 0.98 | (**) | |
| Overall | K | 0.71 | (*) | 0.90 | (**) | 0.93 | (*) | 0.73 | (0.1) | 0.81 | (**) | 1.00 | (**) |
| S | 0.86 | (*) | 0.96 | (**) | 0.95 | (*) | 0.83 | (0.1) | 0.89 | (*) | 1.00 | (**) | |
| P | 0.93 | (**) | 0.98 | (**) | 0.93 | (**) | 0.93 | (**) | 0.90 | (**) | 0.96 | (**) | |
Tab. 1 shows the complete comparison across 2 semantic segmentation tasks, 8 target datasets and 2 microscopy modalities: mitochondria segmentation in 4 EM datasets and nuclei segmentation in 4 LM datasets. The table shows the mean and standard deviation of correlation scores across the set of task-specific target datasets (see Supp. for full tables).
We test CTE with a “hard” (CTE-NHD) and a “soft” (CTE-EI) consistency score, under input (Gauss) and feature space (DropOut) perturbations. In all cases, CTE demonstrates a strong correlation with segmentation performance. The other unsupervised method, Transfer Score (TS), underperforms across all datasets. Supervised metrics fail to take advantage of the target domain labels and do not reliably correlate with the ranking based on F1-score. As they were originally proposed to predict the final fine-tuned performance ranking, we also evaluate them on that task, observing the same lack of correlation as for zero-shot transfer ranking (see Supp.). The only exception is LEEP. It compares source-model predictions with target labels, building an empirical mapping between them and evaluating how well this mapping explains the target data. In direct-transfer settings with shared label sets, LEEP effectively reduces to the average log-likelihood of the correct target class, serving as a proxy for how well the source model’s predictive space separates target labels. Of note, LEEP is also the only baseline that—similar to CTE—operates in the predicted output space and does not make assumptions on the feature space geometry. Interestingly, LEEP is able to correlate for some target datasets (e.g., Hmito), but fails to correlate on average. The key difference between LEEP and LEEP is that the latter fits a Gaussian Mixture Model to the feature space in order to predict pseudo labels; however, it often struggles to converge for the segmentation models.
Tab. 1 shows all CTE variants exhibit strong linear and monotonic correlations with performance across all eight target datasets (four mitochondrial and four nuclei). Comparable results between input- and feature-space perturbations confirm both approaches—previously explored in uncertainty estimation [43, 39], model generalisation [15, 2, 61], and semi-supervised learning [48]—provide effective transferability estimates reflecting the robustness and boundary proximity of target embeddings. For feature-space perturbations, we examined where in the network they should be applied. Prior work [43, 25] suggests intermediate layers, where information is most redundant. To test generality, we applied spatial dropout either to all layers (Mito) or only to the bottleneck (Nuclei). As shown in Tab. 1, both strategies yielded strong correlations. Finally, since our model sets include networks trained both with and without augmentations, the consistently high CTE performance indicates that CTE is robust to training schemes and unaffected by source-domain augmentations during test-time perturbation.
In contrast to the classification case [15], “hard” (CTE-NHD) and “soft” consistency (CTE-EI) metrics perform equivalently well for semantic segmentation.
Uncertainty estimation methods [43] typically quantify output variance across multiple perturbation runs. For transferability ranking, we found that measuring consistency under a single perturbation per test image was sufficient, enabling lightweight model selection with only two inference passes per image. For both input and feature-space perturbations, a broad range of perturbation strengths effectively predicted transfer success via CTE (see Supp.), so all tables show results for single perturbation strengths.
Fig. 3 illustrates CTE performance on a representative target dataset (EPFL mitochondria segmentation). The results show a strong correlation across a diverse set of source models comprising three architectures (U-Net, Residual U-Net, and UNETR), trained on four source datasets (EPFL, Hmito, Rmito, and VNC) with and without input augmentations. All models perform similarly well on their respective source domains, yet visual similarity between source and target images does not reliably predict transferability. Transfer performance is clearly influenced by both architecture and training setup, as seen in the contrast between UNETR (2) and U-Net (3), trained on the same Rmito source data. Our consistency-based approach, by relying on frozen model weights, inherently accounts for both model characteristics and source-domain distributions, enabling a direct assessment of their suitability for a given target dataset.
While CTE generally correlates strongly with transfer performance, pathological cases can cause model mis-rankings. These occur primarily among low-performing transfers, where models exhibit “mode collapse”, producing overly uniform segmentations, such as all background or all foreground. Such degenerate outputs yield artificially high consistency. This limitation could be reduced by pairing consistency with prediction diversity or entropy, which helps distinguish truly stable models from those with collapsed outputs.
| Metric | EPFL | Hmito | Rmito | VNC | ||||||
| pval. | pval. | pval. | pval. | |||||||
| CTE-EI Gauss | K | 0.78 | (**) | 0.75 | (**) | 0.64 | (*) | 0.21 | (0.4) | |
| S | 0.91 | (**) | 0.83 | (**) | 0.81 | (**) | 0.24 | (0.4) | ||
| P | 0.93 | (**) | 0.76 | (**) | 0.70 | (*) | 0.28 | (0.4) | ||
| TS | K | 0.45 | (0.1) | 0.42 | (0.1) | 0.38 | (0.1) | 0.00 | (1.0) | |
| Mean Teacher | S | 0.57 | (0.1) | 0.57 | (0.1) | 0.50 | (0.1) | 0.05 | (0.9) | |
| P | 0.44 | (0.2) | 0.56 | (0.1) | 0.70 | (*) | 0.03 | (0.9) | ||
| CTE-EI Gauss | K | 0.70 | (**) | 0.73 | (**) | 0.55 | (*) | 0.39 | (0.2) | |
| S | 0.86 | (**) | 0.88 | (**) | 0.73 | (**) | 0.47 | (0.2) | ||
| P | 0.76 | (**) | 0.91 | (**) | 0.87 | (**) | 0.71 | (*) | ||
| TS | K | 0.50 | (0.1) | 0.50 | (0.1) | 0.39 | (0.2) | 0.17 | (0.6) | |
| AdaBN | S | 0.67 | (0.1) | 0.73 | (*) | 0.65 | (0.1) | 0.27 | (0.5) | |
| P | 0.90 | (**) | 0.81 | (**) | 0.89 | (**) | 0.34 | (0.4) | ||
When it comes to highly multi-class 3D settings (Tab. 2), both our CTE-EI and CTE-NHD metrics remain reliable, although CTE-NHD is slightly better, probably due to unreliable model calibration in a challenge setting [50]. Correlation scores are computed after aggregating semantically similar classes, i.e., averaging left/right Inferior Alveolar Canals into IACs, the 32 teeth into Teeth, and left/right Maxillary Sinuses into Sinus. We also provide an Overall score, which is the mean across all 42 classes in the ToothFairy2 dataset. Once again, both input and feature space perturbations yield strong and coherent rank correlations. Splitting the class-grouped results reveals our metrics are effective and stable across a morphologically heterogeneous set of classes.
| Metric | Cells | Nuclei | |||||
| K | S | P | K | S | P | ||
| CTE-ARS (Gauss) | 0.69 | 0.83 | 0.90 | 0.72 | 0.83 | 0.79 | |
| ±0.15 | ±0.09 | ±0.04 | ±0.25 | ±0.17 | ±0.21 | ||
| CTE-ARS (DropOut) | 0.69 | 0.83 | 0.91 | 0.80 | 0.85 | 0.84 | |
| ±0.09 | ±0.06 | ±0.04 | ±0.28 | ±0.24 | ±0.11 | ||
Unsupervised Validation. UDA can improve model transfer performance, but if multiple models are adapted, the user again faces the problem of ranking the (fine-tuned) models. Tab. 3 shows the performance of CTE on this task, compared to state-of-the-art Transfer Score. We use mitochondria datasets and two simple UDA methods: a classic Mean Teacher (MT)-based self-supervised domain adaptation with source model predictions as pseudo-labels [65] and adapted batch normalisation (AdaBN) [35]. We observe overall high performance with strong correlation between CTE and final UDA model performance over three out of four target datasets for both UDA approaches. CTE outperforms the previous state-of-the-art for UDA validation, Transfer Score (TS) [78]. The VNC dataset exhibits weaker correlations for both CTE and TS metrics, particularly under the Mean Teacher fine-tuning approach. For CTE, correlation scores are adversely affected by a few outlier models, notably those trained on the EPFL source data. Further inspection revealed these to be pathological transfer cases, where the model consistently segments non-mitochondrial structures in the target data, yielding stable but incorrect predictions and, consequently, low task performance. This misalignment likely arises from distributional differences between the source and target datasets and from the inherent visual similarity of distinct structures in microscopy images, which can induce class confusion during transfer. Consistency regularisation in Mean Teacher training further reinforces this erroneous behaviour. In these cases, the implicit assumption that source and target models address the same task is violated, so models can be ranked well under direct transfer but exhibit outlier behaviour after self-training. The ranking scenario after UDA is inherently more challenging, as the target data effectively becomes part of the model’s adapted source distribution, reducing the impact of perturbations used for CTE estimation. Moreover, the VNC dataset is the smallest of all evaluated targets, containing only 20 slices in total, which further limits statistical robustness and contributes to the weaker correlations observed. As shown in Tab. 3, the AdaBN UDA approach—lacking a consistency component—remains less affected, resulting in higher correlation scores.
5.2 Instance Segmentation
To our knowledge, we are the first to address this problem, so no baseline is available. Unlike the semantic case, instance segmentation does not produce a fixed set of masks with direct correspondence to feature vectors, hence previously compared transferability estimators cannot be applied. For self-trained models we follow a popular instantiation process [53] of predicting cell boundaries and then using seeded watershed and graph partitioning algorithm GASP [6] to obtain instances. Tab. 4 shows the average CTE-ARS performance for instance segmentation, across two tasks and seven target datasets, where it achieves a strong linear and monotonic correlation for all targets.
Evaluating instance model transfer through output prediction consistency enables direct comparison across instance segmentation models employing different instantiation strategies. Fig. 4 presents the transfer ranking of several publicly available models on the Covid-IF cell segmentation dataset: [3] with three transformer backbones (ViT-T, ViT-B, ViT-L), Cellpose-SAM [51], and “Powerful-Chipmunk” [53], a specialist U-Net model trained on this dataset and released via the BioImage Model Zoo [49]. These models differ substantially in architecture and instance formation methodology. The Covid_IF dataset was not included in the training data of , while inclusion for Cellpose-SAM is uncertain. CTE accurately reproduces the true performance ranking across this diverse model set (see Fig. 4).111The ranking does not show the superiority of any model, but is a target dataset specific ranking of automatic segmentation performance.
6 Limitations and Conclusion
The proposed consistency metric assumes source-target task alignment. For generalist models like and Cellpose-SAM, this may not fully hold. The models are finetuned for both cell and nuclei and have no automated mechanism to control which structure is segmented. While supports prompt-based guidance, it requires human input and thus falls outside the scope of automated model ranking. In datasets like Covid_IF, a channel can implicitly contain information about both classes. Models occasionally switch to segmenting nuclei instead of cells (see Supp.). This reduces cell segmentation performance but not consistency scores, as predictions remain internally stable.
Furthermore, the approach can only correlate with pixel-wise metrics that treat all pixels equally, such as Dice or Adapted Rand Error. It is less aligned with boundary-sensitive metrics that assign greater importance to specific regions. In future work, the method could be extended to incorporate pixel-importance weighting within the consistency calculation to address this limitation.
In conclusion, we introduced a novel zero-shot transferability estimator for segmentation models that requires no source training data or target domain labels. Our estimator rankings strongly correlate with direct model performance rankings, achieving state-of-the-art results for semantic and instance segmentation transferability estimation. Our estimators require only two inference runs on test data and apply to blackbox and greybox models, facilitating efficient model selection.
7 Acknowledgements
The research was carried out as part of the AI4Life consortium. AI4Life receives funding from the European Union’s Horizon Europe research and innovation programme under grant agreement number 101057970. Special thanks also to Fynn Beuttenmueller for many constructive conversations.
Supplementary Material
8 Datasets
The general experimental setup for source model transferability ranking is as follows, a set of ‘source’ models are trained for the same task (i.e nuclei semantic segmentation). The models can have a range of architectures and source datasets. A single dataset is then selected as ‘target’ and all of the models from are applied to the test set of the target data. We then rank the transferability of the models and compare this ranking to the transfer performance ranking, obtained using ground-truth. In this section we further detail the employed datasets, and in section Sec. 9 the tested models.
We experimented applying our consistency based transferability estimator (CTE) across a wide range of datasets and modalities from the microscopy and biomedical domains. We focused on ranking the transferability of semantic and instance segmentation models across several distinct tasks. The tasks covered arguably the most common problems in microscopy image analysis; nuclei, cell and mitochondria segmentation. We additionally investigated a popular biomedical segmentation challenge (i.e, ToothFairy2) for human jaw segmentation. Our wide range of experiments across such a diverse set of datasets and tasks highlights the general applicability of our approach for unsupervised ranking of segmentation models.
8.1 Electron Microscopy (EM), Mitochondria
Semantic Segmentation. We use four datasets with semantic segmentation ground-truth, all are neural tissues taken from different species and EM modalities:
EPFL (rat) [40]
The dataset represents a section taken from the CA1 hippocampus region of the brain. The dataset has the rescaled dimensions .
The volume was split with the first 165 z-slices used for training, the next 40 z-slices used for validation and the remaining 125 z-slices used for testing.
MitoEM [17]
The dataset comprises two 3D EM image stacks taken from the brain tissue of a human (Hmito) and a rat (Rmito). Each volume has pixel dimensions .
The volumes were split with the first 300 z-slices used for training, the next 50 z-slices for validation and the remaining 150 z-slices for testing.
VNC (fruit fly) [54]
Transmission Electron Microscopy of ventral nerve cord (VNC) of an adult female Drosophila melanogaster. Once voxels were rescaled the dataset had pixel dimensions .
The VNC dataset is by far the smallest dataset, therefore, in order to maximise available patches for testing we did not perform a typical test train split. For VNC source models we used the first 18 z-slices for training and the remaining 2 z-slices for validation. We did not transfer VNC based models to a VNC test set, removing this transfer from our set of evaluated transfers. For models trained on alternative sources and transferred to VNC we thus tested on the entire 20 z-slices of VNC.
All datasets were resized to match the MitoEM voxel dimensions. For all datasets we split the volumes into 2D image patches for training and evaluation.
Classification. In addition to direct use for semantic segmentation, we reformulated these datasets as a per-patch binary classification task, labelling a patch as positive if mitochondria are present. The volumes were split into 2D patches. To remove positive images with very little mitochondria present, we additionally filtered images based on proportion of mitochondria present in each ground-truth patch. For each dataset we used the corresponding foreground ratios; foreground, foreground, foreground, foreground.
This auxiliary classification task allows us to disentangle the effects of domain shift from the intrinsic suitability of classification-based ranking methods for semantic segmentation.
8.2 Light Microscopy (LM), Nuclei
Semantic Segmentation. We trained nuclei segmentation models on the following fluorescent datasets;
BBBC039 [38]
A high-throughput chemical screen on U2OS cells.
The dataset comprises of a set of 2D images. We used 100 images for training, 50 for validation and 48 for testing. (See github repo for exact image id split).
From each image 2D patches were extracted for training and images were cropped to a standard for evaluation.
Go-Nuclear [68]
Fluorescent dataset of Arabidopsis thaliana.
The dataset comprises of 5 separate data volumes. Three volumes (‘id: 1135’ (), ‘id: 1136’ () and ‘id: 1137’ () were used as training data. For each volume the 40-210 z-slices were taken to avoid imaging artefacts at beginning an end of the volumes. One volume ‘id: 1139’ () was used for validation and the last volume ‘id: 1160’ () was used for testing.
From each volume 2D patches were extracted for training and evaluation.
HeLaCytoNuc [14]
A dataset of HeLa cell nuclei.
The dataset consists of 1873 training images, 535 validation images and 267 test images. Each image has pixel dimensions .
From each image overlapping 2D patches were extracted for training and whole images were taken for evaluation.
Hoechst [5]
A modified U2OS osteosarcoma cell line.
The dataset consists of 30 training images and 10 test images. Each image has pixel dimensions
From each image overlapping 2D patches were extracted for training and whole images were taken for evaluation.
S_BIAD895 [69]
The dataset consists of 47 images with pixel dimensions . We did not split the dataset into test and train splits as we did not apply models trained on S_BIAD895 to S_BIAD895 data, removing this transfer from our set of evaluated transfers. We only considered S_BIAD895 models under transfer to other target datasets.
Thus for models transferred to S_BIAD895, the full image set could be used for testing.
From each image 2D patches were extracted for training and whole images were taken for evaluation.
The SELMA3D 2024 challenge [18].
The dataset consists of eight training volumes, one validation volume and three test volumes.
We took 2D patches from each z-slice for training and evaluation.
S_BIAD1410 [23]
A dataset of fruit fly embryonic development.
The dataset comprises of ten training volumes with the following dimensions ((), (), (), () and six times ()).
From each volume 2D patches were extracted for training and evaluation.
The set of source models trained on the seven available datasets were then transferred to four target datasets; BBBC039 [38], Hoechst [5], S_BIAD895 [69] and DSB2018 [11] (a fluorescent only subset of the data science bowl 2018 challenge). The target datasets were chosen based on two criteria, firstly, for the quality of their ground-truth annotation, to allow for realistic performance evaluation scores. Secondly, the performance of models transferred to the target dataset should have some differentiation. Model differentiation allows for proper investigation of transferability metrics. If all the models transferred to a dataset perform equally then ranking becomes obsolete.
Instance Segmentation. We trained nuclei instance segmentation models on five datasets; BBBC039 [38], HeLaCytoNuc [14], Hoechst [5], S_BIAD895 [69] and SBIAD1410 [23]. The source models were then transferred to the test sets of four target datasets BBBC039 [38], Hoechst [5], S_BIAD895 [69] and one additional dataset S_BIAD634 [31] was introduced. S_BIAD634 is a dataset of nuclei images of normal and cancer cells from different tissue origins and sample preparation types. The dataset comprises of 42 training images and 37 test images. Once again Target datasets were chosen with the same two criteria for instance ground-truth quality and differentiable model performance under transfer.
8.3 Light Microscopy (LM), Cells
Instance Segmentation. We consider four dense instance segmentation problems for cell datasets:
FlyWing [18]
Fluorescent volume of a developing fruit fly wing.
The dataset comprises of four training volumes with the following pixel dimensions ((), (), () and ()), two validation volumes both with dimensions () and two test volumes both with dimensions ().
From each volume 2D patches were extracted for training and full z-slices taken for evaluation.
Ovules [77]
Fluorescent volume of a Arabidopsis thaliana
We used 20 volumes (see data for dimensions) for training, a further single volume () for validation and a single volume () for testing.
From each volume 2D patches were extracted for training and full z-slices taken for evaluation.
PNAS [76]
Fluorescent volume of the Arabidopsis thaliana apical stem cell niche.
For training we used 103 volumes collected at 4 hour time points between 0-84hrs from four plants (see data for dimensions). Volumes from two separate plants were used for validation and testing.
From each volume 2D images were extracted for training and full z-slices taken for evaluation.
Covid-IF [53]
Immunofluorescence imaging of human serum.
The dataset comprises of 48 images, which are all used for testing since we didn’t train any models on this dataset and only use it to evaluate publicly available pre-trained models.
8.4 Cone-Beam Computed Tomography (CBCT), ToothFairy2
Multiclass Semantic Segmentation. For multiclass semantic segmentation experiments, we use ToothFairy2 from a MICCAI2024 challenge [9, 10]. It comprises 530 volumetric scans with voxel-wise annotations of maxillofacial anatomy, of which 480 are publicly available for training, while 50, acquired by a different institution, are held out for evaluation via the grand-challenge platform. The dataset provides annotations for 42 classes, including jawbones, left and right inferior alveolar canals, maxillary sinuses, pharynx, upper and lower teeth, bridges, crowns, and implants. For evaluation, we obtained access to the private test set, collected by a different centre on different acquisition machines.
9 Models
We trained a range of model architectures to build a diverse model set for transfer comparison. All of the models trained by us are trained on the training set of a single ‘source’ dataset. We additionally investigated several publicly available models taken from the Toothfairy2 challenge [9], the Bioimage Model Zoo [49] and large generalist models taken from popular repositories [51, 3]. Information listing the full set of models considered can be found in Sec. 4.2 in the main paper. In the following section we detail the model training setup.
All the self-trained networks were trained until convergence on the source dataset and performed highly on the held out source test set. The classification networks were trained with a Binary Cross entropy (BCE) loss, while the segmentation networks used a combination of Dice and CE losses. For models trained for the Toothfairy2 challenge dataset, the nnUNet planner was used for preprocessing and hyperparameters selection. The models were trained from scratch on a single 48GB A40 Nvidia GPU using CUDA 11.8 and PyTorch 2.1.2. All other models were trained from scratch on a set of 8 12GB NVIDIA GeForce RTX 3080 Ti GPUs using CUDA12.1 and Pytorch 2.5.1.
10 Model Perturbation
Input Perturbations: We tested a range of different input perturbations by applying several popular image transformations at test-time: additive Gaussian noise, gamma correction, and changes in brightness and contrast. The transformations are defined as follows,
where the strength of each perturbation can be controlled by parameters respectively.
Feature Space Perturbations: We relied upon test-time application of spatial DropOut [66] to provide feature space perturbation. We tested several approaches for applying DropOut to our models. Firstly, we tested applying Spatial DropOut only to the bottleneck layer of U-Net style architectures. Our experiments for bottleneck only DropOut are shown on the nuclei datasets both for semantic and instance segmentation (Tabs. 11, 18 and 32). Secondly, in addition to the bottle neck layer we also investigated adding DropOut to all layers that connect the encoder to the decoder, namely skip connections in U-Net style networks. Our experiments for bottleneck + skip connection DropOut are shown on the Toothfairy2 dataset for multiclass semantic segmentation (Tab. 21). Lastly, we investigate applying DropOut uniformly to every layer of the network, this represents the most general and architecture agnostic approach considered. Our experiments for all layer DropOut are shown on the mitochondria semantic segmentation datasets along with the cell instance segmentation datasets (Tabs. 7, 8 and 27).
We found that all the considered feature space perturbation approaches were able to perform effectively, as shown in the tables in the following sections. However, it is worth noting that the first approach of applying DropOut to the bottleneck layer only can be sensitive to residual connections. The second approach of applying DropOut to both the bottleneck layer and skip connections works very well on the model sets tested, but does assume that the models have skip connections in the architecture, otherwise it reduces to the first case (bottleneck only) and could suffer the same sensitivity to residual connections. The last approach of applying DropOut to all layers is the most architecture agnostic and resulted in very strong correlation between the CTE metric and performance score.
11 Semantic Segmentation
In the following section we investigate ranking of semantic segmentation models, each of the tables Tabs. 5, 7, 9, 11, 6, 8, 10 and 18 shows the correlation scores achieved by the CTE metrics for sets of models applied directly to a single target dataset, across a range of perturbation strengths. In the main paper, due to space constraints, we show the correlation scores for a single perturbation strength averaged over the set of target datasets.
We investigate the dependency of CTE on the strength of perturbation applied for both input and feature space perturbations. The tables show that CTE’s performance is stable across a wide range of perturbations.
11.1 Mitochondria Perturbation Sweep
For mitochondria semantic segmentation we investigated applying additive Gaussian noise as an input perturbation between the strengths . We show the results for both CTE-EI (Tab. 5) and CTE-NHD (Tab. 6).
For feature space perturbation we investigated applying spatial DropOut equally to all layers of a network, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for both CTE-EI (Tab. 7) and CTE-NHD (Tab. 8). The tables show that CTE’s performance is stable across a wide range of perturbations.
| Metric | EPFL | Hmito | Rmito | VNC | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI Gauss ( 0.01 - 0.03) | K | 0.73 | (**) | 0.66 | (**) | 0.77 | (**) | 0.46 | (0.06) |
| S | 0.88 | (**) | 0.79 | (**) | 0.90 | (**) | 0.61 | (0.06) | |
| P | 0.82 | (**) | 0.75 | (**) | 0.87 | (**) | 0.63 | (*) | |
| CTE-EI Gauss ( 0.03 - 0.05) | K | 0.77 | (**) | 0.73 | (**) | 0.79 | (**) | 0.52 | (*) |
| S | 0.90 | (**) | 0.86 | (**) | 0.90 | (**) | 0.63 | (*) | |
| P | 0.86 | (**) | 0.82 | (**) | 0.84 | (**) | 0.67 | (*) | |
| CTE-EI Gauss ( 0.05 - 0.07) | K | 0.81 | (**) | 0.73 | (**) | 0.82 | (**) | 0.61 | (**) |
| S | 0.93 | (**) | 0.86 | (**) | 0.92 | (**) | 0.66 | (*) | |
| P | 0.88 | (**) | 0.83 | (**) | 0.82 | (**) | 0.71 | (*) | |
| CTE-EI Gauss ( 0.07 - 0.1) | K | 0.83 | (**) | 0.75 | (**) | 0.90 | (**) | 0.55 | (*) |
| S | 0.93 | (**) | 0.87 | (**) | 0.95 | (**) | 0.63 | (*) | |
| P | 0.91 | (**) | 0.82 | (**) | 0.88 | (**) | 0.73 | (**) | |
| CTE-EI Gauss ( 0.1 - 0.12) | K | 0.87 | (**) | 0.73 | (**) | 0.88 | (**) | 0.58 | (**) |
| S | 0.95 | (**) | 0.87 | (**) | 0.95 | (**) | 0.65 | (*) | |
| P | 0.91 | (**) | 0.80 | (**) | 0.85 | (**) | 0.75 | (**) | |
| CTE-EI Gauss ( 0.12 - 0.15) | K | 0.85 | (**) | 0.73 | (**) | 0.84 | (**) | 0.58 | (**) |
| S | 0.94 | (**) | 0.87 | (**) | 0.94 | (**) | 0.67 | (*) | |
| P | 0.90 | (**) | 0.81 | (**) | 0.84 | (**) | 0.78 | (**) | |
| CTE-EI Gauss ( 0.15 - 0.2) | K | 0.77 | (**) | 0.71 | (**) | 0.84 | (**) | 0.55 | (*) |
| S | 0.92 | (**) | 0.86 | (**) | 0.94 | (**) | 0.62 | (*) | |
| P | 0.86 | (**) | 0.84 | (**) | 0.89 | (**) | 0.78 | (**) | |
| Metric | EPFL | Hmito | Rmito | VNC | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD Gauss ( 0.01 - 0.03) | K | 0.73 | (**) | 0.66 | (**) | 0.71 | (**) | 0.52 | (*) |
| S | 0.89 | (**) | 0.82 | (**) | 0.90 | (**) | 0.63 | (0.05) | |
| P | 0.83 | (**) | 0.75 | (**) | 0.54 | (*) | 0.72 | (**) | |
| CTE-NHD Gauss ( 0.03 - 0.05) | K | 0.77 | (**) | 0.67 | (**) | 0.73 | (**) | 0.48 | (*) |
| S | 0.93 | (**) | 0.83 | (**) | 0.91 | (**) | 0.59 | (0.05) | |
| P | 0.83 | (**) | 0.74 | (**) | 0.61 | (*) | 0.74 | (**) | |
| CTE-NHD Gauss ( 0.05 - 0.07) | K | 0.79 | (**) | 0.68 | (**) | 0.71 | (**) | 0.48 | (*) |
| S | 0.94 | (**) | 0.83 | (**) | 0.90 | (**) | 0.59 | (0.05) | |
| P | 0.84 | (**) | 0.76 | (**) | 0.59 | (*) | 0.69 | (*) | |
| CTE-NHD Gauss ( 0.07 - 0.1) | K | 0.73 | (**) | 0.69 | (**) | 0.75 | (**) | 0.55 | (*) |
| S | 0.91 | (**) | 0.85 | (**) | 0.92 | (**) | 0.64 | (*) | |
| P | 0.90 | (**) | 0.74 | (**) | 0.68 | (**) | 0.72 | (**) | |
| CTE-NHD Gauss ( 0.1 - 0.12) | K | 0.77 | (**) | 0.71 | (**) | 0.77 | (**) | 0.55 | (*) |
| S | 0.92 | (**) | 0.85 | (**) | 0.92 | (**) | 0.64 | (*) | |
| P | 0.92 | (**) | 0.73 | (**) | 0.69 | (**) | 0.75 | (**) | |
| CTE-NHD Gauss ( 0.12 - 0.15) | K | 0.81 | (**) | 0.69 | (**) | 0.77 | (**) | 0.55 | (*) |
| S | 0.93 | (**) | 0.82 | (**) | 0.92 | (**) | 0.64 | (*) | |
| P | 0.91 | (**) | 0.75 | (**) | 0.71 | (**) | 0.77 | (**) | |
| CTE-NHD Gauss ( 0.15 - 0.2) | K | 0.75 | (**) | 0.69 | (**) | 0.75 | (**) | 0.55 | (*) |
| S | 0.91 | (**) | 0.82 | (**) | 0.91 | (**) | 0.65 | (*) | |
| P | 0.88 | (**) | 0.79 | (**) | 0.78 | (**) | 0.77 | (**) | |
| Metric | EPFL | Hmito | Rmito | VNC | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI DropOut ( 0.001) | K | 0.75 | (**) | 0.56 | (**) | 0.73 | (**) | 0.48 | (*) |
| S | 0.93 | (**) | 0.78 | (**) | 0.90 | (**) | 0.60 | (*) | |
| P | 0.91 | (**) | 0.76 | (**) | 0.90 | (**) | 0.78 | (**) | |
| CTE-EI DropOut ( 0.005) | K | 0.75 | (**) | 0.58 | (**) | 0.77 | (**) | 0.48 | (*) |
| S | 0.93 | (**) | 0.79 | (**) | 0.91 | (**) | 0.64 | (*) | |
| P | 0.91 | (**) | 0.76 | (**) | 0.92 | (**) | 0.79 | (**) | |
| CTE-EI DropOut ( 0.01) | K | 0.83 | (**) | 0.58 | (**) | 0.75 | (**) | 0.55 | (*) |
| S | 0.95 | (**) | 0.78 | (**) | 0.90 | (**) | 0.68 | (*) | |
| P | 0.94 | (**) | 0.77 | (**) | 0.92 | (**) | 0.79 | (**) | |
| CTE-EI DropOut ( 0.02) | K | 0.87 | (**) | 0.68 | (**) | 0.77 | (**) | 0.52 | (*) |
| S | 0.96 | (**) | 0.83 | (**) | 0.90 | (**) | 0.66 | (*) | |
| P | 0.95 | (**) | 0.79 | (**) | 0.94 | (**) | 0.79 | (**) | |
| CTE-EI DropOut ( 0.03) | K | 0.85 | (**) | 0.68 | (**) | 0.77 | (**) | 0.61 | (**) |
| S | 0.96 | (**) | 0.83 | (**) | 0.90 | (**) | 0.76 | (**) | |
| P | 0.95 | (**) | 0.79 | (**) | 0.94 | (**) | 0.78 | (**) | |
| CTE-EI DropOut ( 0.04) | K | 0.83 | (**) | 0.66 | (**) | 0.79 | (**) | 0.52 | (*) |
| S | 0.95 | (**) | 0.83 | (**) | 0.91 | (**) | 0.64 | (*) | |
| P | 0.94 | (**) | 0.77 | (**) | 0.91 | (**) | 0.73 | (**) | |
| CTE-EI DropOut ( 0.05) | K | 0.85 | (**) | 0.66 | (**) | 0.79 | (**) | 0.61 | (**) |
| S | 0.96 | (**) | 0.83 | (**) | 0.91 | (**) | 0.73 | (*) | |
| P | 0.95 | (**) | 0.78 | (**) | 0.90 | (**) | 0.71 | (**) | |
| CTE-EI DropOut ( 0.1) | K | 0.75 | (**) | 0.49 | (*) | 0.57 | (**) | 0.52 | (*) |
| S | 0.90 | (**) | 0.65 | (*) | 0.71 | (**) | 0.69 | (**) | |
| P | 0.87 | (**) | 0.68 | (**) | 0.66 | (**) | 0.65 | (*) | |
| Metric | EPFL | Hmito | Rmito | VNC | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD DropOut ( 0.001) | K | 0.69 | (**) | 0.58 | (**) | 0.61 | (**) | 0.58 | (**) |
| S | 0.89 | (**) | 0.76 | (**) | 0.80 | (**) | 0.72 | (*) | |
| P | 0.90 | (**) | 0.69 | (**) | 0.78 | (**) | 0.80 | (**) | |
| CTE-NHD DropOut ( 0.005) | K | 0.71 | (**) | 0.62 | (**) | 0.63 | (**) | 0.61 | (**) |
| S | 0.90 | (**) | 0.80 | (**) | 0.81 | (**) | 0.76 | (**) | |
| P | 0.90 | (**) | 0.70 | (**) | 0.81 | (**) | 0.82 | (**) | |
| CTE-NHD DropOut ( 0.01) | K | 0.77 | (**) | 0.66 | (**) | 0.69 | (**) | 0.64 | (**) |
| S | 0.93 | (**) | 0.82 | (**) | 0.88 | (**) | 0.76 | (*) | |
| P | 0.94 | (**) | 0.72 | (**) | 0.83 | (**) | 0.81 | (**) | |
| CTE-NHD DropOut ( 0.02) | K | 0.81 | (**) | 0.66 | (**) | 0.69 | (**) | 0.61 | (*) |
| S | 0.94 | (**) | 0.81 | (**) | 0.88 | (**) | 0.77 | (**) | |
| P | 0.97 | (**) | 0.73 | (**) | 0.86 | (**) | 0.81 | (**) | |
| CTE-NHD DropOut ( 0.03) | K | 0.81 | (**) | 0.62 | (**) | 0.69 | (**) | 0.64 | (**) |
| S | 0.94 | (**) | 0.78 | (**) | 0.88 | (**) | 0.82 | (**) | |
| P | 0.96 | (**) | 0.74 | (**) | 0.87 | (**) | 0.79 | (**) | |
| CTE-NHD DropOut ( 0.04) | K | 0.81 | (**) | 0.60 | (**) | 0.67 | (**) | 0.48 | (*) |
| S | 0.94 | (**) | 0.76 | (**) | 0.83 | (**) | 0.61 | (0.06) | |
| P | 0.95 | (**) | 0.73 | (**) | 0.85 | (**) | 0.73 | (**) | |
| CTE-NHD DropOut ( 0.05) | K | 0.87 | (**) | 0.62 | (**) | 0.71 | (**) | 0.52 | (*) |
| S | 0.96 | (**) | 0.76 | (**) | 0.84 | (**) | 0.68 | (*) | |
| P | 0.96 | (**) | 0.74 | (**) | 0.85 | (**) | 0.71 | (*) | |
| CTE-NHD DropOut ( 0.1) | K | 0.73 | (**) | 0.47 | (*) | 0.52 | (**) | 0.52 | (*) |
| S | 0.89 | (**) | 0.65 | (**) | 0.65 | (**) | 0.71 | (*) | |
| P | 0.88 | (**) | 0.65 | (**) | 0.64 | (*) | 0.64 | (*) | |
11.2 Nuclei Perturbation Sweep
For nuclei semantic segmentation we investigated applying a range of input augmentations; additive Gaussian noise, Gamma correction, Brightness adjustment and Contrast adjustment (as defined in Sec. 10). We investigated additive Gaussian noise with strengths in the range , Gamma correction with strengths in the range , where equals no adjustment, Brightness adjustment with strengths in the range and Contrast adjustment with strengths in the range , where equals no adjustment. We show the results for both CTE-EI (Tabs. 9, 14, 12 and 13) and CTE-NHD (Tabs. 10, 17, 15 and 16).
For feature space perturbation we investigated applying spatial DropOut only to the bottle neck layer of networks, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for both CTE-EI (Tab. 11) and CTE-NHD (Tab. 18).
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI Gauss ( 0.0 - 0.05) | K | 0.62 | (0.08) | 0.52 | (0.12) | 0.62 | (0.06) | 0.60 | (0.19) |
| S | 0.71 | (0.09) | 0.68 | (0.14) | 0.75 | (0.07) | 0.77 | (0.11) | |
| P | 0.99 | (**) | 0.98 | (**) | 0.95 | (**) | 0.98 | (**) | |
| CTE-EI Gauss ( 0.05 - 0.1) | K | 0.62 | (0.07) | 0.52 | (0.16) | 0.52 | (0.14) | 0.60 | (0.15) |
| S | 0.79 | (*) | 0.68 | (0.11) | 0.71 | (0.10) | 0.77 | (0.10) | |
| P | 1.00 | (**) | 0.99 | (**) | 0.94 | (**) | 0.98 | (**) | |
| CTE-EI Gauss ( 0.1 - 0.2) | K | 0.71 | (*) | 0.52 | (0.16) | 0.52 | (0.13) | 0.73 | (0.06) |
| S | 0.86 | (*) | 0.68 | (0.10) | 0.71 | (0.11) | 0.83 | (0.09) | |
| P | 1.00 | (**) | 0.99 | (**) | 0.93 | (**) | 0.99 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD Gauss ( 0.0 - 0.05) | K | 0.43 | (0.21) | 0.14 | (0.77) | 0.71 | (*) | 0.73 | (0.07) |
| S | 0.64 | (0.12) | 0.18 | (0.69) | 0.86 | (*) | 0.83 | (0.06) | |
| P | 1.00 | (**) | 0.98 | (**) | 0.99 | (**) | 0.99 | (**) | |
| CTE-NHD Gauss ( 0.05 - 0.1) | K | 0.52 | (0.13) | 0.71 | (*) | 0.71 | (*) | 0.47 | (0.27) |
| S | 0.71 | (0.08) | 0.79 | (*) | 0.86 | (*) | 0.54 | (0.30) | |
| P | 1.00 | (**) | 1.00 | (**) | 0.93 | (**) | 0.98 | (**) | |
| CTE-NHD Gauss ( 0.1 - 0.2) | K | 0.71 | (*) | 0.81 | (*) | 0.62 | (0.07) | 0.60 | (0.14) |
| S | 0.86 | (*) | 0.89 | (*) | 0.82 | (*) | 0.71 | (0.12) | |
| P | 1.00 | (**) | 1.00 | (**) | 0.92 | (**) | 0.98 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI DropOut ( 0.05) | K | 0.71 | (*) | 0.43 | (0.24) | 0.52 | (0.18) | 0.60 | (0.13) |
| S | 0.86 | (*) | 0.64 | (0.15) | 0.71 | (0.09) | 0.77 | (0.11) | |
| P | 0.99 | (**) | 0.92 | (**) | 0.90 | (*) | 0.97 | (**) | |
| CTE-EI DropOut ( 0.1) | K | 0.71 | (*) | 0.43 | (0.25) | 0.52 | (0.15) | 0.60 | (0.14) |
| S | 0.86 | (*) | 0.64 | (0.14) | 0.71 | (0.12) | 0.77 | (0.12) | |
| P | 0.99 | (**) | 0.86 | (*) | 0.86 | (*) | 0.96 | (**) | |
| CTE-EI DropOut ( 0.2) | K | 0.71 | (*) | 0.43 | (0.24) | 0.52 | (0.12) | 0.60 | (0.12) |
| S | 0.86 | (*) | 0.64 | (0.13) | 0.71 | (0.09) | 0.77 | (0.08) | |
| P | 0.99 | (**) | 0.86 | (*) | 0.85 | (*) | 0.96 | (**) | |
| CTE-EI DropOut ( 0.3) | K | 0.71 | (*) | 0.52 | (0.12) | 0.52 | (0.14) | 0.60 | (0.15) |
| S | 0.86 | (*) | 0.68 | (0.13) | 0.71 | (0.11) | 0.77 | (0.13) | |
| P | 0.99 | (**) | 0.79 | (*) | 0.85 | (*) | 0.96 | (**) | |
| CTE-EI DropOut ( 0.4) | K | 0.81 | (*) | 0.71 | (*) | 0.52 | (0.15) | 0.73 | (0.06) |
| S | 0.93 | (**) | 0.82 | (*) | 0.71 | (0.09) | 0.83 | (0.06) | |
| P | 0.98 | (**) | 0.83 | (*) | 0.80 | (*) | 0.95 | (**) | |
| CTE-EI DropOut ( 0.5) | K | 0.90 | (*) | 0.81 | (*) | 0.52 | (0.13) | 0.73 | (0.05) |
| S | 0.96 | (*) | 0.89 | (*) | 0.71 | (0.07) | 0.83 | (0.05) | |
| P | 0.98 | (**) | 0.77 | (*) | 0.89 | (*) | 0.95 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI Ctr ( 0.8 - 0.9) | K | 0.81 | (*) | 0.33 | (0.35) | 0.52 | (0.16) | 0.47 | (0.32) |
| S | 0.93 | (*) | 0.50 | (0.26) | 0.68 | (0.12) | 0.71 | (0.11) | |
| P | 0.95 | (**) | 0.91 | (**) | 0.94 | (**) | 0.98 | (**) | |
| CTE-EI Ctr ( 0.9 - 0.95) | K | 0.71 | (*) | 0.33 | (0.36) | 0.52 | (0.15) | 0.47 | (0.26) |
| S | 0.86 | (*) | 0.50 | (0.27) | 0.68 | (0.12) | 0.71 | (0.12) | |
| P | 0.98 | (**) | 0.95 | (**) | 0.95 | (**) | 0.98 | (**) | |
| CTE-EI Ctr ( 1.05 - 1.1) | K | 0.62 | (0.09) | 0.43 | (0.28) | 0.62 | (0.08) | 0.47 | (0.27) |
| S | 0.71 | (0.10) | 0.64 | (0.13) | 0.75 | (0.09) | 0.71 | (0.13) | |
| P | 0.98 | (**) | 0.97 | (**) | 0.94 | (**) | 0.98 | (**) | |
| CTE-EI Ctr ( 1.1 - 1.2) | K | 0.62 | (0.07) | 0.43 | (0.26) | 0.62 | (0.09) | 0.47 | (0.29) |
| S | 0.71 | (0.11) | 0.64 | (0.15) | 0.75 | (0.05) | 0.71 | (0.15) | |
| P | 0.99 | (**) | 0.97 | (**) | 0.95 | (**) | 0.98 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI Gamma ( 0.8 - 0.9) | K | 0.71 | (*) | 0.52 | (0.12) | 0.81 | (**) | 0.60 | (0.12) |
| S | 0.86 | (*) | 0.68 | (0.13) | 0.89 | (*) | 0.77 | (0.09) | |
| P | 0.98 | (**) | 0.96 | (**) | 0.98 | (**) | 0.97 | (**) | |
| CTE-EI Gamma ( 0.9 - 0.95) | K | 0.71 | (*) | 0.52 | (0.14) | 0.71 | (*) | 0.47 | (0.26) |
| S | 0.86 | (*) | 0.68 | (0.13) | 0.82 | (*) | 0.71 | (0.16) | |
| P | 0.99 | (**) | 0.96 | (**) | 0.98 | (**) | 0.96 | (**) | |
| CTE-EI Gamma ( 1.05 - 1.1) | K | 0.71 | (*) | 0.52 | (0.13) | 0.62 | (0.06) | 0.47 | (0.26) |
| S | 0.86 | (*) | 0.68 | (0.12) | 0.75 | (0.08) | 0.71 | (0.14) | |
| P | 0.98 | (**) | 0.94 | (**) | 0.97 | (**) | 0.97 | (**) | |
| CTE-EI Gamma ( 1.1 - 1.2) | K | 0.81 | (*) | 0.52 | (0.16) | 0.81 | (*) | 0.73 | (0.06) |
| S | 0.93 | (**) | 0.68 | (0.14) | 0.89 | (*) | 0.83 | (0.07) | |
| P | 0.96 | (**) | 0.94 | (**) | 0.98 | (**) | 0.96 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI Brt ( 0.0 - 0.05) | K | 0.71 | (*) | 0.43 | (0.25) | 0.52 | (0.16) | 0.60 | (0.12) |
| S | 0.86 | (*) | 0.64 | (0.15) | 0.71 | (0.08) | 0.77 | (0.08) | |
| P | 0.99 | (**) | 0.91 | (**) | 0.94 | (**) | 0.96 | (**) | |
| CTE-EI Brt ( 0.05 - 0.1) | K | 0.81 | (*) | 0.62 | (0.08) | 0.81 | (*) | 0.73 | (0.07) |
| S | 0.93 | (*) | 0.75 | (0.07) | 0.93 | (**) | 0.83 | (0.05) | |
| P | 0.99 | (**) | 0.99 | (**) | 0.94 | (**) | 0.99 | (**) | |
| CTE-EI Brt ( 0.1 - 0.2) | K | 0.71 | (*) | 0.62 | (0.06) | 0.81 | (*) | 0.73 | (0.06) |
| S | 0.86 | (*) | 0.75 | (0.07) | 0.93 | (*) | 0.83 | (0.07) | |
| P | 0.99 | (**) | 0.99 | (**) | 0.94 | (**) | 0.99 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD Ctr ( 0.8 - 0.9) | K | 0.81 | (*) | 0.43 | (0.23) | 0.52 | (0.12) | 0.73 | (0.05) |
| S | 0.93 | (*) | 0.54 | (0.20) | 0.71 | (0.09) | 0.83 | (0.05) | |
| P | 0.94 | (**) | 0.91 | (**) | 0.92 | (**) | 0.99 | (**) | |
| CTE-NHD Ctr ( 0.9 - 0.95) | K | 0.81 | (*) | 0.43 | (0.25) | 0.81 | (*) | 0.60 | (0.12) |
| S | 0.93 | (*) | 0.54 | (0.24) | 0.93 | (*) | 0.77 | (0.09) | |
| P | 0.98 | (**) | 0.96 | (**) | 0.97 | (**) | 0.98 | (**) | |
| CTE-NHD Ctr ( 1.05 - 1.1) | K | 0.81 | (*) | 0.43 | (0.27) | 0.71 | (*) | 0.73 | (0.07) |
| S | 0.93 | (**) | 0.54 | (0.21) | 0.89 | (*) | 0.83 | (0.07) | |
| P | 0.99 | (**) | 0.97 | (**) | 0.92 | (**) | 0.99 | (**) | |
| CTE-NHD Ctr ( 1.1 - 1.2) | K | 0.81 | (*) | 0.43 | (0.24) | 0.71 | (0.05) | 0.60 | (0.14) |
| S | 0.93 | (*) | 0.64 | (0.12) | 0.89 | (*) | 0.77 | (0.08) | |
| P | 0.99 | (**) | 0.99 | (**) | 0.91 | (**) | 0.99 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD Gamma ( 0.8 - 0.9) | K | 0.71 | (*) | 0.62 | (0.05) | 0.90 | (*) | 0.87 | (*) |
| S | 0.86 | (*) | 0.75 | (0.06) | 0.96 | (**) | 0.94 | (*) | |
| P | 0.97 | (**) | 0.91 | (**) | 0.99 | (**) | 0.95 | (**) | |
| CTE-NHD Gamma ( 0.9 - 0.95) | K | 0.62 | (0.07) | 0.52 | (0.14) | 0.81 | (*) | 0.87 | (*) |
| S | 0.75 | (0.06) | 0.68 | (0.13) | 0.93 | (*) | 0.94 | (*) | |
| P | 0.97 | (**) | 0.92 | (**) | 1.00 | (**) | 0.95 | (**) | |
| CTE-NHD Gamma ( 1.05 - 1.1) | K | 0.62 | (0.08) | 0.43 | (0.26) | 1.00 | (**) | 0.73 | (0.05) |
| S | 0.82 | (0.05) | 0.57 | (0.19) | 1.00 | (**) | 0.83 | (0.08) | |
| P | 0.93 | (**) | 0.88 | (*) | 1.00 | (**) | 0.83 | (*) | |
| CTE-NHD Gamma ( 1.1 - 1.2) | K | 0.71 | (*) | 0.62 | (0.06) | 0.81 | (*) | 1.00 | (**) |
| S | 0.79 | (*) | 0.75 | (0.06) | 0.93 | (*) | 1.00 | (*) | |
| P | 0.84 | (*) | 0.89 | (*) | 1.00 | (**) | 0.80 | (0.06) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD Brt ( 0.0 - 0.05) | K | 0.52 | (0.12) | 0.71 | (*) | 0.71 | (*) | 0.47 | (0.25) |
| S | 0.71 | (0.08) | 0.79 | (0.05) | 0.86 | (*) | 0.54 | (0.29) | |
| P | 1.00 | (**) | 0.82 | (*) | 0.92 | (**) | 0.82 | (0.05) | |
| CTE-NHD Brt ( 0.05 - 0.1) | K | 0.71 | (*) | 0.62 | (0.07) | 0.71 | (*) | 0.73 | (0.05) |
| S | 0.86 | (*) | 0.71 | (0.08) | 0.86 | (*) | 0.83 | (0.06) | |
| P | 1.00 | (**) | 0.98 | (**) | 0.90 | (*) | 0.99 | (**) | |
| CTE-NHD Brt ( 0.1 - 0.2) | K | 0.71 | (*) | 0.71 | (*) | 0.90 | (*) | 0.73 | (0.07) |
| S | 0.86 | (*) | 0.82 | (*) | 0.96 | (**) | 0.83 | (0.07) | |
| P | 1.00 | (**) | 0.99 | (**) | 0.93 | (**) | 0.99 | (**) | |
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-NHD DropOut ( 0.05) | K | 0.71 | (*) | 0.43 | (0.23) | 0.52 | (0.15) | 0.87 | (*) |
| S | 0.86 | (*) | 0.64 | (0.16) | 0.68 | (0.11) | 0.94 | (*) | |
| P | 0.97 | (**) | 0.01 | (0.98) | 0.65 | (0.11) | 0.81 | (0.05) | |
| CTE-NHD DropOut ( 0.1) | K | 0.71 | (*) | 0.71 | (0.05) | 0.62 | (0.06) | 0.73 | (0.05) |
| S | 0.86 | (*) | 0.86 | (*) | 0.82 | (*) | 0.83 | (0.07) | |
| P | 0.97 | (**) | -0.03 | (0.95) | 0.61 | (0.15) | 0.68 | (0.13) | |
| CTE-NHD DropOut ( 0.2) | K | 0.90 | (*) | 0.52 | (0.12) | 0.71 | (*) | 0.73 | (0.07) |
| S | 0.96 | (*) | 0.68 | (0.13) | 0.89 | (*) | 0.83 | (0.05) | |
| P | 0.97 | (**) | 0.10 | (0.82) | 0.67 | (0.10) | 0.74 | (0.10) | |
| CTE-NHD DropOut ( 0.3) | K | 0.90 | (*) | 0.62 | (0.07) | 0.81 | (*) | 0.73 | (0.05) |
| S | 0.96 | (**) | 0.71 | (0.10) | 0.93 | (*) | 0.83 | (0.06) | |
| P | 0.97 | (**) | -0.01 | (0.99) | 0.63 | (0.13) | 0.74 | (0.09) | |
| CTE-NHD DropOut ( 0.4) | K | 0.81 | (*) | 0.71 | (0.05) | 0.71 | (*) | 0.60 | (0.13) |
| S | 0.89 | (*) | 0.89 | (*) | 0.89 | (*) | 0.71 | (0.14) | |
| P | 0.97 | (**) | 0.18 | (0.70) | 0.63 | (0.13) | 0.71 | (0.11) | |
| CTE-NHD DropOut ( 0.5) | K | 0.81 | (*) | 0.71 | (*) | 0.52 | (0.13) | 0.73 | (0.05) |
| S | 0.89 | (*) | 0.89 | (*) | 0.68 | (0.13) | 0.83 | (0.05) | |
| P | 0.96 | (**) | 0.11 | (0.81) | 0.84 | (*) | 0.68 | (0.14) | |
| Metric | EPFL | Hmito | Rmito | VNC | |||||
| pval. | pval. | pval. | pval. | ||||||
| CCFV | K | 0.03 | (0.89) | -0.09 | (0.70) | -0.21 | (0.30) | -0.18 | (0.45) |
| S | 0.02 | (0.96) | -0.15 | (0.56) | -0.29 | (0.30) | -0.22 | (0.51) | |
| P | -0.03 | (0.92) | -0.17 | (0.54) | -0.23 | (0.40) | -0.29 | (0.36) | |
| NLEEP | K | 0.26 | (0.21) | 0.77 | (**) | 0.65 | (**) | -0.06 | (0.86) |
| S | 0.32 | (0.23) | 0.90 | (**) | 0.82 | (**) | -0.06 | (0.92) | |
| P | 0.26 | (0.34) | 0.66 | (*) | 0.63 | (*) | -0.01 | (0.97) | |
| GBC | K | 0.45 | (*) | 0.43 | (*) | 0.34 | (0.09) | 0.33 | (0.15) |
| S | 0.60 | (*) | 0.54 | (*) | 0.51 | (0.06) | 0.41 | (0.17) | |
| P | 0.65 | (*) | 0.54 | (*) | 0.47 | (0.08) | 0.23 | (0.48) | |
| Hscore | K | -0.12 | (0.58) | 0.33 | (0.10) | 0.36 | (0.06) | 0.36 | (0.11) |
| S | -0.21 | (0.46) | 0.50 | (0.05) | 0.53 | (*) | 0.50 | (0.09) | |
| P | 0.00 | (1.00) | 0.49 | (0.07) | 0.46 | (0.08) | 0.54 | (0.07) | |
| LEEP | K | 0.79 | (**) | 0.90 | (**) | 0.94 | (**) | 0.91 | (**) |
| S | 0.89 | (**) | 0.98 | (**) | 0.98 | (**) | 0.98 | (**) | |
| P | 0.96 | (**) | 0.96 | (**) | 0.98 | (**) | 0.98 | (**) | |
| LogME | K | -0.24 | (0.23) | 0.10 | (0.61) | 0.25 | (0.22) | 0.21 | (0.39) |
| S | -0.33 | (0.23) | 0.17 | (0.56) | 0.32 | (0.24) | 0.25 | (0.43) | |
| P | -0.30 | (0.27) | 0.05 | (0.85) | 0.09 | (0.75) | 0.41 | (0.18) | |
| NCTI | K | -0.05 | (0.85) | 0.30 | (0.14) | 0.27 | (0.16) | 0.33 | (0.13) |
| S | -0.07 | (0.75) | 0.42 | (0.12) | 0.35 | (0.23) | 0.44 | (0.17) | |
| P | 0.07 | (0.81) | 0.42 | (0.12) | 0.42 | (0.12) | 0.58 | (0.05) | |
| RegHscore | K | -0.12 | (0.52) | 0.31 | (0.12) | 0.40 | (*) | 0.36 | (0.10) |
| S | -0.21 | (0.44) | 0.49 | (0.06) | 0.58 | (*) | 0.50 | (0.11) | |
| P | 0.00 | (1.00) | 0.48 | (0.07) | 0.46 | (0.08) | 0.54 | (0.07) | |
| TS | K | 0.54 | (*) | 0.28 | (0.14) | 0.08 | (0.77) | 0.09 | (0.74) |
| S | 0.72 | (*) | 0.33 | (0.23) | 0.02 | (0.94) | 0.11 | (0.70) | |
| P | 0.60 | (*) | 0.19 | (0.50) | 0.14 | (0.62) | 0.00 | (0.99) | |
11.3 Baseline Transferability Metrics
We also show the correlation scores per dataset for all the baseline transferability metrics. Tab. 19 shows the results for the four mitochondria datasets (EPFL, Hmito, Rmito and VNC) and Tab. 20 shows the results for the four nuclei datasets (BBBC039, DSB2018, Hoechst, S_BIAD895). In the main paper due to space constraints we showed the correlation scores averaged within the dataset groups.
| Metric | BBBC039 | DSB2018 | Hoechst | S_BIAD895 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CCFV | K | 0.05 | (1.00) | 0.05 | (1.00) | 0.24 | (0.54) | -0.47 | (0.29) |
| S | 0.14 | (0.77) | 0.04 | (0.98) | 0.21 | (0.61) | -0.60 | (0.25) | |
| P | -0.31 | (0.51) | -0.01 | (0.99) | 0.45 | (0.31) | -0.78 | (0.07) | |
| GBC | K | -0.14 | (0.80) | -0.14 | (0.82) | 0.24 | (0.54) | -0.47 | (0.29) |
| S | -0.21 | (0.67) | -0.21 | (0.65) | 0.36 | (0.45) | -0.60 | (0.21) | |
| P | -0.16 | (0.72) | 0.11 | (0.81) | 0.32 | (0.48) | -0.59 | (0.22) | |
| LEEP | K | 0.81 | (*) | 0.71 | (*) | 0.14 | (0.77) | 1.00 | (**) |
| S | 0.89 | (*) | 0.86 | (*) | 0.32 | (0.51) | 1.00 | (*) | |
| P | 0.95 | (**) | 0.92 | (**) | 0.94 | (**) | 0.94 | (*) | |
| NLEEP | K | 0.14 | (0.77) | -0.05 | (1.00) | 0.14 | (0.77) | 0.20 | (0.79) |
| S | 0.18 | (0.72) | -0.04 | (1.00) | 0.14 | (0.76) | 0.20 | (0.70) | |
| P | 0.62 | (0.14) | 0.30 | (0.52) | 0.09 | (0.85) | 0.26 | (0.62) | |
| Hscore | K | 0.33 | (0.33) | -0.14 | (0.78) | -0.05 | (1.00) | 1.00 | (*) |
| S | 0.36 | (0.46) | -0.36 | (0.45) | 0.00 | (1.00) | 1.00 | (*) | |
| P | 0.36 | (0.43) | 0.13 | (0.77) | -0.04 | (0.94) | 0.78 | (0.07) | |
| RegHscore | K | 0.33 | (0.41) | 0.05 | (0.97) | 0.14 | (0.72) | 0.87 | (*) |
| S | 0.36 | (0.45) | -0.07 | (0.92) | 0.21 | (0.66) | 0.94 | (*) | |
| P | 0.29 | (0.52) | 0.18 | (0.70) | 0.07 | (0.88) | 0.84 | (*) | |
| LogME | K | 0.14 | (0.82) | -0.14 | (0.78) | -0.33 | (0.41) | -0.47 | (0.25) |
| S | 0.18 | (0.67) | -0.21 | (0.68) | -0.50 | (0.24) | -0.60 | (0.23) | |
| P | -0.36 | (0.43) | -0.54 | (0.21) | -0.12 | (0.80) | -0.94 | (*) | |
| NCTI | K | 0.33 | (0.39) | -0.14 | (0.76) | -0.05 | (1.00) | 1.00 | (**) |
| S | 0.36 | (0.45) | -0.36 | (0.45) | 0.00 | (1.00) | 1.00 | (*) | |
| P | 0.37 | (0.41) | 0.14 | (0.76) | -0.02 | (0.96) | 0.72 | (0.10) | |
| TS | K | 0.14 | (0.84) | -0.33 | (0.41) | -0.05 | (1.00) | 0.33 | (0.45) |
| S | 0.25 | (0.64) | -0.43 | (0.35) | 0.11 | (0.77) | 0.49 | (0.35) | |
| P | -0.02 | (0.97) | -0.56 | (0.20) | -0.13 | (0.79) | 0.37 | (0.48) | |
11.4 Multiclass Toothfairy2 Perturbation Sweep
For multiclass semantic segmentation we investigated applying both additive Gaussian noise (controlled by ) and Gamma Correction (controlled by ) as input perturbations. For feature space perturbation we investigated applying spatial DropOut to the bottleneck and skip connection layers of the network, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We report the performance of both CTE-EI (Tab. 21) and CTE-NHD (Tab. 22).
| Metric | IACs | Teeth | Mand. | Sinus. | Overall | ||||||
| pval. | pval. | pval. | pval. | pval. | |||||||
| CTE-EI Gauss ( 2.0x) | K | 0.90 | (**) | 0.81 | (**) | 0.62 | (*) | 0.43 | (0.3) | 0.71 | (*) |
| S | 0.96 | (**) | 0.89 | (*) | 0.82 | (*) | 0.50 | (0.3) | 0.86 | (*) | |
| P | 0.99 | (**) | 0.95 | (**) | 0.73 | (0.1) | 0.38 | (0.4) | 0.93 | (**) | |
| CTE-EI Gauss ( 2.5x) | K | 0.90 | (**) | 0.81 | (**) | 0.62 | (0.1) | 0.43 | (0.3) | 0.71 | (*) |
| S | 0.96 | (**) | 0.89 | (*) | 0.82 | (*) | 0.50 | (0.3) | 0.86 | (*) | |
| P | 0.99 | (**) | 0.94 | (**) | 0.70 | (0.1) | 0.38 | (0.4) | 0.93 | (**) | |
| CTE-EI Gauss ( 3.0x) | K | 0.90 | (**) | 0.81 | (*) | 0.62 | (0.1) | 0.43 | (0.2) | 0.71 | (*) |
| S | 0.96 | (**) | 0.89 | (**) | 0.82 | (*) | 0.50 | (0.3) | 0.86 | (*) | |
| P | 0.99 | (**) | 0.92 | (**) | 0.63 | (0.1) | 0.39 | (0.4) | 0.92 | (**) | |
| CTE-EI Gamma ( 0.2) | K | 0.90 | (**) | 0.71 | (*) | 0.62 | (0.1) | 0.52 | (0.2) | 0.81 | (**) |
| S | 0.96 | (**) | 0.86 | (*) | 0.75 | (0.1) | 0.57 | (0.2) | 0.89 | (*) | |
| P | 0.99 | (**) | 0.68 | (0.1) | 0.61 | (0.2) | 0.46 | (0.3) | 0.73 | (0.1) | |
| CTE-EI Gamma ( 0.5) | K | 1.00 | (**) | 0.81 | (**) | 0.81 | (*) | 0.43 | (0.3) | 0.81 | (**) |
| S | 1.00 | (**) | 0.93 | (**) | 0.89 | (**) | 0.43 | (0.3) | 0.89 | (**) | |
| P | 0.96 | (**) | 0.91 | (**) | 0.81 | (*) | 0.44 | (0.3) | 0.91 | (**) | |
| CTE-EI Gamma ( 2.0) | K | 1.00 | (**) | 0.71 | (*) | 0.43 | (0.3) | 0.43 | (0.2) | 0.81 | (**) |
| S | 1.00 | (**) | 0.86 | (*) | 0.54 | (0.2) | 0.57 | (0.2) | 0.89 | (*) | |
| P | 0.99 | (**) | 0.89 | (**) | 0.74 | (0.1) | 0.38 | (0.4) | 0.90 | (**) | |
| CTE-EI DropOut ( 0.3) | K | 1.00 | (**) | 0.52 | (0.2) | 0.81 | (*) | 0.37 | (0.1) | 0.62 | (0.1) |
| S | 1.00 | (**) | 0.75 | (0.1) | 0.82 | (*) | 0.51 | (0.2) | 0.79 | (*) | |
| P | 0.93 | (**) | 0.92 | (**) | 0.89 | (**) | 0.40 | (0.2) | 0.90 | (**) | |
| CTE-EI DropOut ( 0.5) | K | 0.87 | (*) | 0.67 | (0.1) | 0.87 | (*) | 0.87 | (*) | 0.93 | (*) |
| S | 0.94 | (**) | 0.77 | (0.1) | 0.87 | (*) | 0.84 | (*) | 0.95 | (*) | |
| P | 0.83 | (*) | 0.85 | (**) | 0.89 | (**) | 0.88 | (**) | 0.93 | (**) | |
| CTE-EI DropOut ( 0.7) | K | 0.71 | (*) | 0.55 | (0.3) | 0.71 | (*) | 0.53 | (0.3) | 0.54 | (0.2) |
| S | 0.86 | (*) | 0.60 | (0.4) | 0.82 | (0.1) | 0.51 | (0.2) | 0.42 | (0.4) | |
| P | 0.73 | (0.1) | 0.82 | (*) | 0.89 | (**) | 0.50 | (0.3) | 0.79 | (*) | |
| Metric | IACs | Teeth | Mand. | Sinus. | Overall | ||||||
| pval. | pval. | pval. | pval. | pval. | |||||||
| CTE-NHD Gauss ( 2.0x) | K | 0.81 | (**) | 1.00 | (**) | 0.52 | (0.2) | 0.81 | (*) | 0.90 | (**) |
| S | 0.93 | (*) | 1.00 | (**) | 0.68 | (0.1) | 0.89 | (*) | 0.96 | (**) | |
| P | 0.98 | (**) | 0.98 | (**) | 0.65 | (0.1) | 0.73 | (0.1) | 0.98 | (**) | |
| CTE-NHD Gauss ( 2.5x) | K | 0.81 | (**) | 1.00 | (**) | 0.43 | (0.2) | 0.81 | (**) | 0.90 | (**) |
| S | 0.93 | (**) | 1.00 | (**) | 0.57 | (0.2) | 0.89 | (*) | 0.96 | (**) | |
| P | 0.99 | (**) | 0.96 | (**) | 0.60 | (0.2) | 0.67 | (**) | 0.96 | (**) | |
| CTE-NHD Gauss ( 3.0x) | K | 0.81 | (**) | 1.00 | (**) | 0.52 | (0.2) | 0.81 | (**) | 0.90 | (**) |
| S | 0.93 | (**) | 1.00 | (**) | 0.68 | (0.1) | 0.89 | (*) | 0.96 | (**) | |
| P | 0.99 | (**) | 0.94 | (**) | 0.50 | (0.3) | 0.64 | (0.1) | 0.94 | (**) | |
| CTE-NHD Gamma ( 0.2) | K | 0.81 | (*) | 0.71 | (*) | 0.43 | (0.3) | 0.62 | (0.1) | 0.81 | (**) |
| S | 0.93 | (**) | 0.86 | (*) | 0.49 | (0.1) | 0.75 | (0.1) | 0.89 | (**) | |
| P | 0.99 | (**) | 0.67 | (0.1) | 0.42 | (0.2) | 0.64 | (0.1) | 0.72 | (0.1) | |
| CTE-NHD Gamma ( 0.5) | K | 0.90 | (**) | 0.81 | (*) | 0.90 | (**) | 0.81 | (*) | 0.81 | (**) |
| S | 0.96 | (**) | 0.93 | (**) | 0.96 | (**) | 0.89 | (**) | 0.89 | (**) | |
| P | 0.96 | (**) | 0.94 | (**) | 0.79 | (*) | 0.97 | (**) | 0.94 | (**) | |
| CTE-NHD Gamma ( 2.0) | K | 0.90 | (*) | 0.90 | (**) | 0.81 | (**) | 0.90 | (**) | 1.00 | (**) |
| S | 0.96 | (**) | 0.96 | (**) | 0.89 | (**) | 0.96 | (**) | 1.00 | (**) | |
| P | 0.98 | (**) | 0.95 | (**) | 0.80 | (*) | 0.98 | (**) | 0.96 | (**) | |
| CTE-NHD DropOut ( 0.3) | K | 0.90 | (**) | 0.71 | (*) | 0.81 | (**) | 0.43 | (0.2) | 0.81 | (**) |
| S | 0.96 | (**) | 0.86 | (*) | 0.89 | (*) | 0.43 | (0.4) | 0.89 | (**) | |
| P | 0.86 | (**) | 0.95 | (**) | 0.83 | (*) | 0.77 | (*) | 0.95 | (**) | |
| CTE-NHD DropOut ( 0.5) | K | 0.87 | (*) | 0.75 | (0.1) | 0.87 | (**) | 0.83 | (*) | 0.73 | (0.1) |
| S | 0.94 | (**) | 0.77 | (0.1) | 0.94 | (*) | 0.94 | (*) | 0.83 | (0.1) | |
| P | 0.80 | (0.1) | 0.94 | (**) | 0.96 | (**) | 0.90 | (**) | 0.93 | (**) | |
| CTE-NHD DropOut ( 0.7) | K | 0.62 | (0.1) | 0.43 | (0.2) | 0.81 | (*) | 0.44 | (0.2) | 0.53 | (0.3) |
| S | 0.71 | (0.1) | 0.50 | (0.3) | 0.93 | (**) | 0.43 | (0.3) | 0.49 | (0.2) | |
| P | 0.69 | (0.1) | 0.85 | (*) | 0.86 | (**) | 0.57 | (0.2) | 0.71 | (*) | |
11.5 Post-UDA Perturbation Sweep
In this section we investigate the performance of the CTE metric for predicting the model transferability ranking after the application of two unsupervised domain adaptation (UDA) approaches, Mean Teacher [65] (Tab. 23) and adaptive batch normalisation (AdaBN) [35] (Tab. 24). The tables show a sweep over a range of perturbation strengths and report the correlation scores per target dataset. In the main paper due to space limitations we reported the correlation scores averaged over the target datasets for a single perturbation strength. We investigated applying additive Gaussian noise to the input of the models with strengths varying in the range .
| Metric | EPFL | Hmito | Rmito | VNC | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI Gauss ( 0.01 - 0.03) | K | 0.71 | (**) | 0.64 | (**) | 0.56 | (*) | 0.15 | (0.57) |
| S | 0.85 | (**) | 0.78 | (*) | 0.75 | (*) | 0.22 | (0.53) | |
| P | 0.93 | (**) | 0.71 | (*) | 0.52 | (0.10) | 0.11 | (0.73) | |
| CTE-EI Gauss ( 0.03 - 0.05) | K | 0.82 | (**) | 0.67 | (**) | 0.60 | (*) | 0.09 | (0.71) |
| S | 0.90 | (**) | 0.79 | (**) | 0.78 | (*) | 0.15 | (0.65) | |
| P | 0.95 | (**) | 0.75 | (**) | 0.59 | (0.05) | 0.16 | (0.61) | |
| CTE-EI Gauss ( 0.05 - 0.07) | K | 0.82 | (**) | 0.71 | (**) | 0.64 | (**) | 0.09 | (0.75) |
| S | 0.90 | (**) | 0.81 | (**) | 0.80 | (*) | 0.17 | (0.61) | |
| P | 0.96 | (**) | 0.76 | (**) | 0.64 | (*) | 0.19 | (0.55) | |
| CTE-EI Gauss ( 0.07 - 0.1) | K | 0.78 | (**) | 0.67 | (**) | 0.67 | (**) | 0.12 | (0.65) |
| S | 0.89 | (**) | 0.78 | (*) | 0.85 | (**) | 0.16 | (0.63) | |
| P | 0.96 | (**) | 0.77 | (**) | 0.67 | (*) | 0.22 | (0.49) | |
| CTE-EI Gauss ( 0.1 - 0.12) | K | 0.82 | (**) | 0.75 | (**) | 0.64 | (*) | 0.12 | (0.60) |
| S | 0.92 | (**) | 0.83 | (**) | 0.81 | (**) | 0.16 | (0.65) | |
| P | 0.96 | (**) | 0.77 | (**) | 0.69 | (*) | 0.27 | (0.39) | |
| CTE-EI Gauss ( 0.12 - 0.15) | K | 0.78 | (**) | 0.75 | (**) | 0.64 | (*) | 0.21 | (0.38) |
| S | 0.91 | (**) | 0.83 | (**) | 0.81 | (**) | 0.24 | (0.45) | |
| P | 0.93 | (**) | 0.76 | (**) | 0.70 | (*) | 0.28 | (0.38) | |
| CTE-EI Gauss ( 0.15 - 0.2) | K | 0.67 | (**) | 0.71 | (**) | 0.64 | (*) | 0.18 | (0.45) |
| S | 0.85 | (**) | 0.82 | (**) | 0.81 | (**) | 0.22 | (0.47) | |
| P | 0.89 | (**) | 0.72 | (*) | 0.71 | (*) | 0.28 | (0.37) | |
| P | 0.72 | (*) | 0.65 | (*) | 0.70 | (*) | 0.26 | (0.41) | |
| Metric | EPFL | Hmito | Rmito | VNC | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-EI Gauss ( 0.01 - 0.03) | K | 0.67 | (**) | 0.73 | (**) | 0.55 | (*) | 0.33 | (0.30) |
| S | 0.80 | (**) | 0.85 | (**) | 0.71 | (*) | 0.45 | (0.25) | |
| P | 0.93 | (**) | 0.93 | (**) | 0.92 | (**) | 0.57 | (0.11) | |
| CTE-EI Gauss ( 0.03 - 0.05) | K | 0.70 | (**) | 0.73 | (**) | 0.61 | (**) | 0.33 | (0.21) |
| S | 0.87 | (**) | 0.88 | (**) | 0.76 | (*) | 0.42 | (0.29) | |
| P | 0.89 | (**) | 0.92 | (**) | 0.89 | (**) | 0.62 | (0.07) | |
| CTE-EI Gauss ( 0.05 - 0.07) | K | 0.73 | (**) | 0.73 | (**) | 0.58 | (*) | 0.33 | (0.25) |
| S | 0.88 | (**) | 0.88 | (**) | 0.74 | (**) | 0.42 | (0.26) | |
| P | 0.85 | (**) | 0.91 | (**) | 0.88 | (**) | 0.71 | (*) | |
| CTE-EI Gauss ( 0.07 - 0.1) | K | 0.70 | (**) | 0.73 | (**) | 0.55 | (*) | 0.39 | (0.16) |
| S | 0.86 | (**) | 0.88 | (**) | 0.73 | (*) | 0.47 | (0.21) | |
| P | 0.76 | (**) | 0.91 | (**) | 0.87 | (**) | 0.71 | (*) | |
| CTE-EI Gauss ( 0.1 - 0.12) | K | 0.67 | (**) | 0.73 | (**) | 0.58 | (**) | 0.33 | (0.28) |
| S | 0.83 | (**) | 0.88 | (**) | 0.76 | (**) | 0.42 | (0.27) | |
| P | 0.59 | (*) | 0.90 | (**) | 0.86 | (**) | 0.70 | (*) | |
| CTE-EI Gauss ( 0.12 - 0.15) | K | 0.67 | (**) | 0.73 | (**) | 0.61 | (**) | 0.39 | (0.16) |
| S | 0.81 | (**) | 0.90 | (**) | 0.77 | (**) | 0.52 | (0.15) | |
| P | 0.44 | (0.15) | 0.89 | (**) | 0.85 | (**) | 0.71 | (*) | |
| CTE-EI Gauss ( 0.15 - 0.2) | K | 0.64 | (**) | 0.67 | (**) | 0.58 | (*) | 0.39 | (0.17) |
| S | 0.76 | (*) | 0.85 | (**) | 0.76 | (**) | 0.52 | (0.15) | |
| P | 0.38 | (0.22) | 0.86 | (**) | 0.82 | (**) | 0.74 | (*) | |
11.6 Final-Finetuned Correlation Transfer Metrics
For the supervised baseline methods compared against [47, 7, 79, 52, 27, 36, 74, 80] the goal of the transferability metrics is to predict the final supervised finetuned transfer performance of a model based on its initial pre-trained state. In the main paper we investigate the real world scenario of unsupervised model transfer, hence we are limited to assessing direct zero-shot application of pre-trained models to target data or assessing models after application of UDA approaches. We showed already in Tab. 1 that, even with supervision, existing transfer metrics fail at direct zero-shot transferability estimation in the semantic setting. How about in the setting of predicting the final supervised finetuned performance of models?
To investigate this we further trained a set of source models on the supervised training set of each of the mitochondria datasets in turn and then used the available ground truth to evaluate the performance of the fine-tuned models on the test set. Then using the pre-finetuning weights of the models we calculated all the base line transferability metrics and measured the correlation scores between pre-finetuned transferability metric and post-finetuned performance score. Tab. 25 shows the correlation scores split by target dataset, we can see that all the metrics fail to meaningfully correlate with final-finetuned performance. Further highlighting that existing transferability metrics cannot be reliably used for predicting the transfer of semantic segmentation models applied to biomedical data.
| Metric | EPFL | Hmito | Rmito | VNC | |||||
| pval. | pval. | pval. | pval. | ||||||
| CCFV | K | -0.16 | (0.57) | -0.42 | (0.10) | -0.53 | (*) | -0.27 | (0.25) |
| S | -0.24 | (0.48) | -0.57 | (0.05) | -0.69 | (*) | -0.37 | (0.24) | |
| P | -0.25 | (0.46) | -0.52 | (0.10) | -0.50 | (0.12) | -0.40 | (0.19) | |
| TS | K | 0.31 | (0.24) | 0.49 | (*) | 0.38 | (0.11) | -0.06 | (0.88) |
| S | 0.55 | (0.09) | 0.63 | (0.05) | 0.56 | (0.07) | -0.04 | (0.87) | |
| P | 0.20 | (0.56) | 0.69 | (*) | 0.68 | (*) | -0.16 | (0.62) | |
| LEEP | K | 0.56 | (*) | -0.09 | (0.77) | 0.16 | (0.55) | 0.27 | (0.28) |
| S | 0.72 | (*) | -0.04 | (0.92) | 0.36 | (0.31) | 0.41 | (0.20) | |
| P | 0.44 | (0.17) | 0.10 | (0.78) | 0.39 | (0.23) | 0.37 | (0.23) | |
| NCTI | K | -0.16 | (0.54) | -0.09 | (0.75) | 0.16 | (0.57) | 0.00 | (1.00) |
| S | -0.21 | (0.54) | -0.24 | (0.51) | 0.26 | (0.45) | -0.01 | (0.99) | |
| P | 0.08 | (0.82) | -0.24 | (0.49) | 0.37 | (0.26) | 0.04 | (0.89) | |
| LogME | K | -0.13 | (0.59) | -0.02 | (0.98) | 0.35 | (0.17) | 0.00 | (1.00) |
| S | 0.05 | (0.84) | -0.01 | (0.97) | 0.41 | (0.22) | 0.13 | (0.66) | |
| P | 0.16 | (0.64) | -0.08 | (0.82) | 0.03 | (0.94) | 0.29 | (0.36) | |
| Hscore | K | -0.02 | (1.00) | 0.27 | (0.26) | 0.45 | (0.06) | 0.15 | (0.49) |
| S | 0.09 | (0.81) | 0.42 | (0.24) | 0.53 | (0.12) | 0.28 | (0.35) | |
| P | 0.20 | (0.56) | 0.21 | (0.54) | 0.41 | (0.20) | 0.17 | (0.60) | |
| GBC | K | 0.31 | (0.21) | 0.20 | (0.47) | 0.45 | (0.07) | 0.24 | (0.34) |
| S | 0.47 | (0.14) | 0.25 | (0.45) | 0.48 | (0.16) | 0.28 | (0.36) | |
| P | 0.24 | (0.47) | 0.53 | (0.10) | 0.41 | (0.22) | 0.45 | (0.14) | |
| RegHscore | K | -0.02 | (1.00) | 0.24 | (0.37) | 0.45 | (0.06) | 0.15 | (0.58) |
| S | 0.09 | (0.78) | 0.40 | (0.22) | 0.53 | (0.09) | 0.28 | (0.40) | |
| P | 0.20 | (0.56) | 0.21 | (0.54) | 0.41 | (0.21) | 0.17 | (0.60) | |
| NLEEP | K | 0.09 | (0.78) | -0.20 | (0.41) | 0.38 | (0.09) | 0.03 | (0.92) |
| S | 0.15 | (0.66) | -0.21 | (0.52) | 0.45 | (0.17) | -0.03 | (0.92) | |
| P | 0.11 | (0.75) | 0.10 | (0.76) | 0.45 | (0.17) | -0.02 | (0.95) | |
12 Instance Segmentation
In the following section we investigate ranking of instance segmentation models, each of the tables Tabs. 26, 27, 28, 29, 31, 30, 32 and 33 shows the correlation scores achieved by the CTE-ARS metric for sets of models applied directly to a single target dataset, across a range of perturbation strengths. In the main paper, due to space constraints, we show the correlation scores for a single perturbation strength averaged over the set of target datasets. We investigate the dependency of CTE on the strength of perturbation applied for both input and feature space perturbations. The tables show that CTE’s performance is stable across a wide range of perturbations.
12.1 Cells Perturbation Sweep
For cell instance segmentation we investigated applying additive Gaussian noise as an input perturbation between the strengths . We show the results for CTE-ARS in Tab. 26.
For feature space perturbation we investigated applying spatial DropOut equally to all layers of a network, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for both CTE-ARS in Tab. 27.
| Metric | FlyWing | Ovules | PNAS | ||||
| pval. | pval. | pval. | |||||
| CTE-ARS Gauss ( 0.01 - 0.05) | K | 0.64 | (*) | 0.43 | (0.16) | 0.93 | (**) |
| S | 0.79 | (*) | 0.60 | (0.16) | 0.98 | (**) | |
| P | 0.79 | (*) | 0.95 | (**) | 0.83 | (*) | |
| CTE-ARS Gauss ( 0.05 - 0.1) | K | 0.57 | (0.07) | 0.43 | (0.21) | 0.71 | (*) |
| S | 0.76 | (*) | 0.57 | (0.15) | 0.83 | (*) | |
| P | 0.91 | (**) | 0.87 | (**) | 0.91 | (**) | |
| CTE-ARS Gauss ( 0.1 - 0.15) | K | 0.57 | (0.07) | 0.64 | (*) | 0.57 | (0.07) |
| S | 0.76 | (*) | 0.79 | (*) | 0.76 | (*) | |
| P | 0.94 | (**) | 0.88 | (**) | 0.89 | (**) | |
| CTE-ARS Gauss ( 0.15 - 0.2) | K | 0.50 | (0.11) | 0.57 | (0.06) | 0.57 | (0.06) |
| S | 0.74 | (0.05) | 0.74 | (0.05) | 0.76 | (*) | |
| P | 0.94 | (**) | 0.85 | (*) | 0.90 | (**) | |
| CTE-ARS Gauss ( 0.2 - 0.25) | K | 0.57 | (0.07) | 0.64 | (*) | 0.86 | (**) |
| S | 0.76 | (*) | 0.79 | (*) | 0.93 | (**) | |
| P | 0.94 | (**) | 0.88 | (**) | 0.87 | (**) | |
| Metric | FlyWing | Ovules | PNAS | ||||
| pval. | pval. | pval. | |||||
| CTE-ARS DropOut ( 0.001) | K | 0.64 | (*) | 0.50 | (0.11) | 0.71 | (*) |
| S | 0.81 | (*) | 0.74 | (0.06) | 0.83 | (*) | |
| P | 0.93 | (**) | 0.82 | (*) | 0.81 | (*) | |
| CTE-ARS DropOut ( 0.005) | K | 0.71 | (*) | 0.50 | (0.14) | 0.71 | (*) |
| S | 0.83 | (*) | 0.74 | (*) | 0.83 | (*) | |
| P | 0.93 | (**) | 0.83 | (*) | 0.86 | (*) | |
| CTE-ARS DropOut ( 0.01) | K | 0.57 | (0.07) | 0.50 | (0.11) | 0.71 | (*) |
| S | 0.76 | (*) | 0.74 | (0.06) | 0.83 | (*) | |
| P | 0.94 | (**) | 0.84 | (*) | 0.88 | (**) | |
| CTE-ARS DropOut ( 0.02) | K | 0.64 | (*) | 0.64 | (*) | 0.79 | (*) |
| S | 0.79 | (*) | 0.79 | (*) | 0.90 | (*) | |
| P | 0.94 | (**) | 0.87 | (**) | 0.92 | (**) | |
| CTE-ARS DropOut ( 0.03) | K | 0.50 | (0.11) | 0.79 | (*) | 0.64 | (*) |
| S | 0.74 | (0.07) | 0.90 | (*) | 0.81 | (*) | |
| P | 0.95 | (**) | 0.89 | (**) | 0.91 | (**) | |
| CTE-ARS DropOut ( 0.04) | K | 0.50 | (0.12) | 0.57 | (0.07) | 0.64 | (*) |
| S | 0.74 | (*) | 0.76 | (*) | 0.81 | (*) | |
| P | 0.95 | (**) | 0.79 | (*) | 0.86 | (*) | |
| CTE-ARS DropOut ( 0.05) | K | 0.50 | (0.10) | 0.57 | (0.05) | 0.71 | (*) |
| S | 0.74 | (0.06) | 0.71 | (0.05) | 0.83 | (*) | |
| P | 0.95 | (**) | 0.83 | (*) | 0.88 | (**) | |
| CTE-ARS DropOut ( 0.1) | K | 0.57 | (0.06) | 0.43 | (0.17) | 0.57 | (*) |
| S | 0.76 | (*) | 0.60 | (0.14) | 0.71 | (0.07) | |
| P | 0.95 | (**) | 0.78 | (*) | 0.88 | (**) | |
12.2 Nuclei Perturbation Sweep
For nuclei instance segmentation we investigated applying a range of input augmentations; additive Gaussian noise, Gamma correction, Brightness adjustment and Contrast adjustment (as defined in Sec. 10). We investigated additive Gaussian noise with strengths in the range , Gamma correction with strengths in the range , where equals no adjustment, Brightness adjustment with strengths in the range and Contrast adjustment with strengths in the range , where equals no adjustment. We show the results for both CTE-ARS (Tabs. 28, 29, 30 and 31)
For feature space perturbation we investigated applying spatial DropOut only to the bottle neck layer of networks, where the strength of the perturbation is controlled by the proportion of feature maps dropped at each layer . We investigated DropOut proportions in the range . We show results for for CTE-ARS in Tab. 32.
We again observe stable correlation performance across a wide range of perturbation strengths. Although, as discussed in the main paper, Sec. 3.4, care should be taken to ensure perturbations remain ‘tolerable’ in the extreme perturbation case, too strong of a perturbation can lead to a loss of correlation between consistency based metrics and transfer performance. For example, the last row of Tab. 32 shows the strongest level of DropOut () applied can cause the CTE-ARS metric to become unstable on S_BIAD634.
| Metric | BBBC039 | Hoechst | S_BIAD895 | S_BIAD634 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-ARS Gauss ( 0.0 - 0.05) | K | 0.80 | (0.08) | 0.80 | (0.11) | 0.33 | (0.75) | 0.80 | (0.10) |
| S | 0.90 | (0.10) | 0.90 | (0.08) | 0.60 | (0.42) | 0.90 | (0.08) | |
| P | 0.99 | (**) | 0.85 | (0.07) | 0.47 | (0.53) | 0.93 | (*) | |
| CTE-ARS Gauss ( 0.05 - 0.1) | K | 0.20 | (0.80) | 0.80 | (0.09) | 0.33 | (0.75) | 1.00 | (*) |
| S | 0.50 | (0.42) | 0.90 | (0.08) | 0.40 | (0.75) | 1.00 | (*) | |
| P | 0.72 | (0.17) | 0.89 | (*) | 0.99 | (*) | 0.80 | (0.10) | |
| CTE-ARS Gauss ( 0.1 - 0.2) | K | 0.40 | (0.54) | 1.00 | (*) | 0.67 | (0.33) | 0.80 | (0.09) |
| S | 0.60 | (0.36) | 1.00 | (*) | 0.80 | (0.33) | 0.90 | (0.08) | |
| P | 0.50 | (0.39) | 0.93 | (*) | 0.94 | (0.06) | 0.79 | (0.11) | |
| Metric | BBBC039 | Hoechst | S_BIAD895 | S_BIAD634 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-ARS Gamma ( 0.8 - 0.9) | K | 0.40 | (0.42) | 0.80 | (0.08) | -0.33 | (0.75) | 0.40 | (0.49) |
| S | 0.60 | (0.37) | 0.90 | (0.10) | -0.40 | (0.75) | 0.60 | (0.36) | |
| P | 0.62 | (0.26) | 0.82 | (0.09) | 0.06 | (0.94) | 0.64 | (0.25) | |
| CTE-ARS Gamma ( 0.9 - 0.95) | K | 0.40 | (0.50) | 0.80 | (0.08) | 0.33 | (0.75) | 0.80 | (0.08) |
| S | 0.60 | (0.37) | 0.90 | (0.10) | 0.40 | (0.75) | 0.90 | (0.07) | |
| P | 0.54 | (0.35) | 0.89 | (*) | 0.58 | (0.42) | 0.76 | (0.14) | |
| CTE-ARS Gamma ( 1.05 - 1.1) | K | 0.60 | (0.22) | 0.80 | (0.09) | 0.00 | (1.00) | 0.80 | (0.08) |
| S | 0.70 | (0.22) | 0.90 | (0.08) | -0.20 | (0.92) | 0.90 | (0.09) | |
| P | 0.94 | (*) | 0.84 | (0.07) | 0.44 | (0.56) | 0.82 | (0.09) | |
| CTE-ARS Gamma ( 1.1 - 1.2) | K | 0.40 | (0.53) | 1.00 | (*) | 0.33 | (0.75) | 0.40 | (0.45) |
| S | 0.60 | (0.37) | 1.00 | (*) | 0.40 | (0.75) | 0.60 | (0.34) | |
| P | 0.71 | (0.18) | 0.87 | (0.05) | 0.68 | (0.32) | 0.76 | (0.14) | |
| Metric | BBBC039 | Hoechst | S_BIAD895 | S_BIAD634 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-ARS Ctr ( 0.8 - 0.9) | K | 0.80 | (0.09) | 1.00 | (*) | 0.00 | (1.00) | 0.80 | (0.08) |
| S | 0.90 | (0.09) | 1.00 | (*) | -0.20 | (0.92) | 0.90 | (0.07) | |
| P | 0.85 | (0.07) | 0.94 | (*) | 0.18 | (0.82) | 0.83 | (0.08) | |
| CTE-ARS Ctr ( 0.9 - 0.95) | K | 0.80 | (0.09) | 1.00 | (*) | 0.67 | (0.33) | 0.80 | (0.10) |
| S | 0.90 | (0.09) | 1.00 | (*) | 0.80 | (0.33) | 0.90 | (0.09) | |
| P | 0.96 | (*) | 0.99 | (**) | 0.52 | (0.48) | 0.87 | (0.05) | |
| CTE-ARS Ctr ( 1.05 - 1.1) | K | 0.74 | (0.13) | 1.00 | (*) | 0.67 | (0.33) | 0.60 | (0.25) |
| S | 0.82 | (0.13) | 1.00 | (*) | 0.80 | (0.33) | 0.80 | (0.13) | |
| P | 0.99 | (**) | 0.95 | (*) | 0.58 | (0.42) | 0.80 | (0.10) | |
| CTE-ARS Ctr ( 1.1 - 1.2) | K | 0.80 | (0.11) | 0.60 | (0.23) | 0.33 | (0.75) | 0.80 | (0.07) |
| S | 0.90 | (0.10) | 0.70 | (0.20) | 0.40 | (0.75) | 0.90 | (0.09) | |
| P | 0.83 | (0.08) | 0.81 | (0.10) | 0.38 | (0.62) | 0.90 | (*) | |
| Metric | BBBC039 | Hoechst | S_BIAD895 | S_BIAD634 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-ARS Brt ( 0.0 - 0.05) | K | 0.60 | (0.22) | 0.60 | (0.26) | 0.00 | (1.00) | 0.80 | (0.10) |
| S | 0.70 | (0.25) | 0.70 | (0.21) | 0.20 | (0.92) | 0.90 | (0.09) | |
| P | 0.80 | (0.10) | 0.83 | (0.08) | 0.77 | (0.23) | 0.72 | (0.17) | |
| CTE-ARS Brt ( 0.05 - 0.1) | K | 0.40 | (0.46) | 0.80 | (0.11) | 0.33 | (0.75) | 0.80 | (0.07) |
| S | 0.60 | (0.32) | 0.90 | (0.10) | 0.40 | (0.75) | 0.90 | (0.10) | |
| P | 0.69 | (0.19) | 0.92 | (*) | 0.80 | (0.20) | 0.84 | (0.08) | |
| CTE-ARS Brt ( 0.1 - 0.2) | K | 0.60 | (0.23) | 0.80 | (0.09) | 0.67 | (0.33) | 0.40 | (0.49) |
| S | 0.70 | (0.23) | 0.90 | (0.10) | 0.80 | (0.33) | 0.50 | (0.40) | |
| P | 0.81 | (0.10) | 0.85 | (0.07) | 0.99 | (*) | 0.43 | (0.47) | |
| Metric | BBBC039 | Hoechst | S_BIAD895 | S_BIAD634 | |||||
| pval. | pval. | pval. | pval. | ||||||
| CTE-ARS DropOut ( 0.05) | K | 0.40 | (0.43) | 0.80 | (0.08) | 1.00 | (0.08) | 0.80 | (0.12) |
| S | 0.60 | (0.32) | 0.90 | (0.08) | 1.00 | (0.08) | 0.90 | (0.08) | |
| P | 0.83 | (0.08) | 0.90 | (*) | 0.93 | (0.07) | 0.88 | (0.05) | |
| CTE-ARS DropOut ( 0.1) | K | 0.40 | (0.51) | 1.00 | (*) | 1.00 | (0.08) | 0.80 | (0.08) |
| S | 0.50 | (0.44) | 1.00 | (*) | 1.00 | (0.08) | 0.90 | (0.11) | |
| P | 0.87 | (0.05) | 0.86 | (0.06) | 0.93 | (0.07) | 0.91 | (*) | |
| CTE-ARS DropOut ( 0.2) | K | 0.40 | (0.54) | 1.00 | (*) | 1.00 | (0.08) | 0.80 | (0.09) |
| S | 0.50 | (0.45) | 1.00 | (*) | 1.00 | (0.08) | 0.90 | (0.07) | |
| P | 0.95 | (*) | 0.93 | (*) | 0.72 | (0.28) | 0.78 | (0.12) | |
| CTE-ARS DropOut ( 0.3) | K | 0.60 | (0.25) | 1.00 | (*) | 1.00 | (0.08) | 0.80 | (0.08) |
| S | 0.70 | (0.23) | 1.00 | (*) | 1.00 | (0.08) | 0.90 | (0.10) | |
| P | 0.87 | (0.05) | 0.89 | (*) | 0.76 | (0.24) | 0.82 | (0.09) | |
| CTE-ARS DropOut ( 0.4) | K | 0.60 | (0.22) | 1.00 | (*) | 0.67 | (0.33) | 1.00 | (*) |
| S | 0.70 | (0.24) | 1.00 | (*) | 0.80 | (0.33) | 1.00 | (*) | |
| P | 0.95 | (*) | 0.91 | (*) | 0.91 | (0.09) | 0.96 | (*) | |
| CTE-ARS DropOut ( 0.5) | K | 0.60 | (0.22) | 0.80 | (0.10) | 0.67 | (0.33) | 0.00 | (1.00) |
| S | 0.70 | (0.25) | 0.90 | (0.07) | 0.80 | (0.33) | 0.00 | (1.00) | |
| P | 0.98 | (**) | 0.80 | (0.10) | 0.89 | (0.11) | 0.49 | (0.40) | |
12.3 Covid_IF External Models Perturbation Sweep
In the main paper in Fig. 4 we also ranked the transferability of a set of publicly available models [51, 3, 53] for instance cell segmentation on the Covid_IF [53] dataset. The models all have very different instantiation processes, but through calculating the transferability metric in output segmentation space we are able to directly compare the varied set of models. The Micro-SAM models [3] perform Automatic Instance Segmentation (AIS), which utilises an additionally trained decoder to predict a three channel output: the distance to the object centre, the distance to the object boundary and foreground probabilities. The predictions are then instantiated using seeded watershed. Cellpose-SAM [51] use a custom decoder to directly predict vector flow fields, which can then be converted to instance segmentations. ‘Powerful-chipmunk’ is a specialist U-Net model taken from the Bioimage model zoo [49], the model is trained specifically on the Covid_IF data and predicts object foreground and boundary predictions, which are then converted to instances using watershed and GASP.
In Fig. 4 from the main paper we show the correlation between CTE-ARS and mSA performance for a single additive Gaussian noise perturbation strength. In Tab. 33 we investigate a range of Gaussian noise strengths .
| Metric | Covid_IF | ||
| pval. | |||
| CTE-ARS Gauss ( 0.01 - 0.02) | K | 0.80 | (0.09) |
| S | 0.90 | (0.08) | |
| P | 0.86 | (0.06) | |
| CTE-ARS Gauss ( 0.02 - 0.03) | K | 1.00 | (*) |
| S | 1.00 | (*) | |
| P | 0.93 | (*) | |
| CTE-ARS Gauss ( 0.03 - 0.04) | K | 1.00 | (*) |
| S | 1.00 | (*) | |
| P | 0.91 | (*) | |
| CTE-ARS Gauss ( 0.04 - 0.05) | K | 1.00 | (*) |
| S | 1.00 | (*) | |
| P | 0.93 | (*) | |
12.4 Cells vs Nuclei Prediction
A key assumption of the proposed consistency based approach is that the source and target tasks align, i.e, if a model was trained for nuclei segmentation then it should only be evaluated for nuclei segmentation. For generalist models like [3] and Cellpose-SAM [51], this assumption may not fully hold. The models are finetuned for both cell and nuclei and have no automated mechanism to control which structure is segmented. In datasets like Covid_IF, a channel can implicitly contain information about both classes. We observed that Cellpose-SAM and occasionally seemingly switch to segmenting nuclei instead of cells (Figs. 5 and 6) particularly in ‘challenging’ sections of the image. It is hard to quantify is this always due to a ‘mode’ switch to nuclei segmentation or just poor cell segmentation performance, however it seems clear that at least in some cases the models segment small objects (much smaller than cells) that clearly correspond to the nuclei information present in the serum channel. This ‘mode’ switch reduces cell segmentation performance, but not consistency scores, as predictions remain internally stable.